### v1

# Tapping into the Black Box: Uncovering Aligned Representations in Pretrained Neural Networks

Abstract
In this paper we argue that ReLU networks learn an implicit linear model we can actually tap into. We describe that alleged model formally and show that we can approximately pull its decision boundary back to the input space with certain simple modification to the backward pass. The resulting gradients (called excitation pullbacks) reveal high-resolution input- and target-specific features of remarkable perceptual alignment on a number of popular ImageNet-pretrained deep architectures. This strongly suggests that neural networks do, in fact, rely on learned interpretable patterns that can be recovered after training. Thus, our findings may have profound implications for knowledge discovery and the development of dependable artificial systems.
Figure 1: A rudimentary 5-step pixel-space gradient ascent guided by excitation pullbacks for pretrained ResNet50. Each cell shows the difference between the perturbed and clean image, targeting the class in the column. Diagonal: original class; off-diagonal: counterfactuals. Last column: randomly selected extra label. See Section 5 for details.
1 Introduction
This work investigates gradient-based explanations in ReLU networks, with a focus on convolutional architectures commonly used in computer vision. While our empirical results and theoretical developments concentrate on standard ReLU networks trained on image classification tasks, the core arguments are sufficiently general to apply - with the appropriate modifications - to other activation functions and modalities.
We introduce a simple but effective method for interpreting neural predictions via excitation pullbacks, which isolate the robust informative signal in neural activations from the noise introduced by local data peculiarities. We motivate this idea extensively and demonstrate its empirical effectiveness on the following ImageNet-pretrained ReLU models: ResNet50 (He et al., 2015), VGG11_BN (Simonyan and Zisserman, 2015), and DenseNet121 (Huang et al., 2018). These specific architectures were chosen for reasons detailed in Section 5. We also give compelling theoretical arguments backing the claim that excitation pullbacks are indeed faithful local approximations of the network’s decision boundary. In particular, we hypothesize that they reveal a simpler linear model approximated by the trained network, encoded in the pre-activation patterns.
The paper outline is as follows:
  • Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks are described;
  • Section 3 defines tensor field pullbacks and shows that neural networks are linear in their path space;
  • Section 4 defines excitation pullbacks and postulates that they approximate the decision boundary of an implicit kernel machine that is learned in the network’s pre-activation patterns;
  • Section 5 shows that excitation pullbacks of selected architectures pretrained on the ImageNet are perceptually aligned and can be used to produce faithful and meaningful local feature accentuations by just a few steps of gradient ascent along their directions;
  • Section 6 describes the related works, indicates a possible biological feasibility of presented methods and discusses their limitations.
This paper thus bridges practical gradient visualization with theoretical insights into neural function, suggesting a path toward more transparent and robust deep models.
2 Preliminaries and notational conventions
This section establishes important notions and conventions used throughout the rest of the paper.
2.1 ReLU networks
Let $M$ be a feedforward ReLU network with parameters $\theta\in\mathbb{R}^{d_{net}}$, expressible as a composition of affine transformations and pointwise ReLU nonlinearities.111we omit max-pooling layers from the analysis to improve clarity, while noting that similar arguments are likely to apply due to the structural similarity between max-pooling and ReLU; see Section 5.3. Let $L$ denote the number of layers, and let $d_{\ell}$ be the width of the $\ell$-th layer, with $d_{0}$ denoting the input dimension. Let $x_{0}$ denote the identity function on $\mathbb{R}^{d_{0}}$, i.e. $x_{0}(x):=x$. The network applies the following operations:
$z_{\ell}=W_{\ell}x_{\ell-1}+b_{\ell},\quad x_{\ell}=\text{ReLU}(z_{\ell}),\qquad\text{for }\ell=1,\dotsc,L$
where $W_{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$, $b_{\ell}\in\mathbb{R}^{d_{\ell}}$, and $\text{ReLU}(z)=\max(0,z)$ is applied element-wise. We define the overall network function $M:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ by the recursive application of the layers, i.e.,
$M(x_{0}):=x_{L}$
In this paper we’ll treat $M$ as backbone for some linear model, i.e. $x_{L}$ represents a penultimate layer in some chosen architecture. Specifically, we’ll consider $f:\mathbb{R}^{d_{0}}\to\mathbb{R}$ defined by:
$f_{c}:=y_{c}^{\top}M$
where $y_{c}$ is a trainable set of weights $y_{c}\in\mathbb{R}^{d_{L}}$ (a neuron in the network’s head) and $y_{c}^{\top}$ is it’s corresponding covector, i.e. the linear action of $y_{c}$ on $\mathbb{R}^{d_{L}}$ by the dot product: $y_{c}^{\top}x=\langle y_{c},x\rangle$. Standard classification network with $C$ classes has $C$ corresponding vectors $y_{c}$. We will skip the lower index $c$ assuming that it’s fixed.
For notational brevity, we consider augmented matrices and vectors where the bias term is incorporated by appending a 1 to the end of the input vector. Specifically, we define the augmented input and weight matrix as
$\tilde{x}_{\ell-1}=\begin{bmatrix}x_{\ell-1}\\ 1\end{bmatrix}\in\mathbb{R}^{d_{\ell-1}+1},\quad\tilde{W}_{\ell}=\begin{bmatrix}W_{\ell}&b_{\ell}\\ 0\quad\cdots\quad 0&1\end{bmatrix}\in\mathbb{R}^{(d_{\ell}+1)\times(d_{\ell-1}+1)}$
Then, the augmented pre-activation vector is given by
$\tilde{z}_{\ell}=\tilde{W}_{\ell}\tilde{x}_{\ell-1}$
Conventions
Throughout the paper, we will often abuse notation by omitting the tilde and the explicit increase of dimension, writing simply
$z_{\ell}=W_{\ell}x_{\ell-1}$
with the understanding that the bias is included in the matrix $W_{\ell}$ and the input vector $x_{\ell-1}$. We will occasionally write $\mathbb{R}^{d^{*}_{0}}\subset\mathbb{R}^{d_{0}}$ to denote the true input space (without the implicitly added coordinate). Notice that, by construction, the $\mathbb{R}^{d^{*}_{0}}$ remains the actual domain of $f$.
2.2 Gating representation of ReLU networks
In ReLU networks, each activation can be equivalently viewed as a gating mechanism applied to the pre-activations. Specifically, for every layer $\ell\in\{1,\dots,L\}$, the activation function
$x_{\ell}=\text{ReLU}(z_{\ell})$
can be rewritten as
$x_{\ell}=g_{\ell}\odot z_{\ell}$
where $g_{\ell}\in\{0,1\}^{d_{\ell}}$ is a binary gating vector defined by
$g_{\ell}[i]=\begin{cases}1&\text{if }z_{\ell}[i]>0,\\ 0&\text{otherwise}.\end{cases}$
Since $z_{\ell}$ is itself a function of the input $x_{0}$, we may define a gating function
$G:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\{0,1\}^{d_{\ell}},\quad G(x):=(g_{1},\dots,g_{L})$ (1)
Let $G_{\ell}(x)=\mathrm{diag}(g_{\ell})\in\mathbb{R}^{d_{\ell}\times d_{\ell}}$ denote the diagonal matrix whose diagonal entries are given by $g_{\ell}$. Then, the network output $M(x)=x_{L}$ can be written as a product of (augmented) weight matrices interleaved with input-dependent diagonal matrices:
$M(x)=G_{L}(x)W_{L}G_{L-1}(x)W_{L-1}\cdots G_{1}(x)W_{1}x$
where each $W_{\ell}$ includes the bias terms via the convention introduced earlier.
This representation makes the piecewise-linear structure of ReLU networks explicit: for a fixed input $x$, the gating pattern $(g_{1},\dots,g_{L})\in\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ determines a purely linear computation.
Similarly, let’s define the pre-activation function as:
$Z:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}},\quad Z(x):=(z_{1},\dots,z_{L})$ (2)
i.e. the concatenation of all pre-activations of $M(x)$.
2.3 Gating-induced networks
More generally, given any gating function:
$\Lambda:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}},\quad\Lambda:=(\lambda_{1},\dots,\lambda_{L}),\quad\lambda_{\ell}:\mathbb{R}^{d_{0}}\to[0,1]^{d_{\ell}}$ (3)
we can define the $\Lambda$-induced network as
$M_{\Lambda}(x):=\Lambda_{L}(x)W_{L}\Lambda_{L-1}(x)W_{L-1}\cdots\Lambda_{1}(x)W_{1}x$ (4)
where $\Lambda_{\ell}(x)=\mathrm{diag}(\lambda_{\ell}(x))\in\mathbb{R}^{d_{\ell}\times d_{\ell}}$ is the diagonal matrix formed from the $\ell$-th component $\lambda_{\ell}$ of $\Lambda(x)$. In particular, $M$ is $G$-induced, i.e. $M=M_{G}$. Note that $M_{\Lambda}$ implicitly depends on the network parameters $\theta$.
We can also define the matrix field $\vec{M}_{\Lambda}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}\times\mathbb{R}^{d_{0}}$ by:
$\vec{M}_{\Lambda}(x):=\Lambda_{L}(x)W_{L}\Lambda_{L-1}(x)W_{L-1}\cdots\Lambda_{1}(x)W_{1}$ (5)
so that $M_{\Lambda}(x)=\vec{M}_{\Lambda}(x)x$. We set $f_{\Lambda}:=y^{\top}M_{\Lambda}$ and therefore $f=f_{G}$.
2.4 Pullbacks and vector fields
Note: the symbols in this subsection 2.4 are unrelated to the rest of the section 2.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a scalar-valued function, and let $\phi:\mathbb{R}^{d_{0}}\to\mathbb{R}^{n}$ be a map. The pullback of $f$ through $\phi$, denoted $\phi^{*}f$, is defined as the composition
$\phi^{*}f:=f\circ\phi$
which is a function $\phi^{*}f:\mathbb{R}^{d_{0}}\to\mathbb{R}$ given explicitly by
$(\phi^{*}f)(x)=f(\phi(x))$
This construction allows us to view $f$ as a scalar function defined directly on the domain $\mathbb{R}^{d_{0}}$ (through the action of $\phi$).
Convention
Whenever a piecewise continuous vector field $v:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{0}}$ satisfies
$\langle v(x),x\rangle=v^{\top}(x)x=f(\phi(x))=(\phi^{*}f)(x)$
we will refer both to $v$ and it’s corresponding covector field $v^{\top}:\mathbb{R}^{d_{0}}\to(\mathbb{R}^{d_{0}})^{*}$ as pullbacks (of $f$ through $\phi$).
2.5 Gating-induced pullbacks
We can pull the linear map $y^{\top}:\mathbb{R}^{d_{L}}\to\mathbb{R}$ back to the input space $\mathbb{R}^{d_{0}}$ through $M_{\Lambda}$, generating the following covector field over $\mathbb{R}^{d_{0}}$:
$v^{\top}_{\Lambda}:=y^{\top}\vec{M}_{\Lambda}$
We will call both $v^{\top}_{\Lambda}$ and the corresponding vector field $v_{\Lambda}$ a $\Lambda$-pullback. In particular, we have:
$v_{G}^{\top}=y^{\top}\vec{M}_{G}$ (6)
Now, the following sequence of equations can be easily verified:
$f_{\Lambda}(x)=y^{\top}M_{\Lambda}(x)=y^{\top}\vec{M}_{\Lambda}(x)x=v_{\Lambda}^{\top}(x)x=\langle v_{\Lambda}(x),x\rangle$ (7)
This means that $f_{\Lambda}$ is represented by it’s $\Lambda$-pullback $v_{\Lambda}$.
Now, since $f$ is locally affine in $\mathbb{R}^{d^{*}_{0}}$ (see Section 2.1), then:
$\nabla f=\nabla f_{G}=(y^{\top}\vec{M}_{G})|_{\mathbb{R}^{d^{*}_{0}}}=(v_{G}^{\top})|_{\mathbb{R}^{d^{*}_{0}}}$ (8)
This means that $v_{G}|_{\mathbb{R}^{d_{0}^{*}}}$ can be easily computed as the network’s gradient.
Note 2.1 The pullback representation is important for explainability as it expresses the network action as a dot product with a well-understood entity, i.e. an input-space vector (field).
2.6 Training dynamics
[…]
3 Neural networks are linear in their path space
3.1 Network action as atom filtering
Let’s dive deeper into the structure of $M$. Fix a hidden neuron $u_{\ell}[i]$, i.e. the one corresponding to the $i$-th output in the $\ell$-th layer of $M$. Due to the compositional nature of the network, we can define the pullback $v_{\Lambda}|_{(\ell,i)}$ of $u_{\ell}[i]$ in the same way we did for $y$ just consider the appropriate subnetwork below $u_{\ell}[i]$ as the backbone. Moreover, the full-network pullback $v_{\Lambda}$ can be expressed as a weighted sum of the pullbacks of neurons from previous layers, for example:
$v_{\Lambda}=\sum_{i=1}^{d_{L}}y[i]\cdot\lambda_{L}[i]\cdot v_{\Lambda}|_{(L,i)}$
[…]
We can also express $v_{\Lambda}$ as the weighted sum of first-layer pullbacks:
$v_{\Lambda}=\sum_{(i_{1},i_{2},\ldots,i_{L})}y[i_{L}]\cdot\lambda_{L}[i_{L}]\cdot W_{L}[i_{L},i_{L-1}]\cdots\lambda_{2}[i_{2}]\cdot W_{2}[i_{2},i_{1}]\cdot\lambda_{1}[i_{1}]\cdot v_{\Lambda}|_{(1,i_{1})}$
Where $(i_{1},\ldots,i_{L})$ is the path through layers $1,\ldots,L$, i.e. the selection of indices $i_{\ell}\in\{1,\ldots,d_{\ell}\}$. Now rearrange the factors:
$v_{\Lambda}=\sum_{(i_{1},i_{2},\ldots,i_{L})}\left(\prod_{\ell=1}^{L}\lambda_{\ell}[i_{\ell}]\right)\left(y[i_{L}]\prod_{l=2}^{L}W_{\ell}[i_{\ell},i_{{\ell-1}}]\right)\cdot v_{\Lambda}|_{(1,i_{1})}$
Notice that $v_{G}|_{(1,i_{1})}=(W_{1}[i_{1}])^{\top}\in\mathbb{R}^{d_{0}}$, i.e. it’s a $i_{1}$-th row of $W_{1}$, so we have:
$v_{\Lambda}=\sum_{(i_{1},i_{2},\ldots,i_{L})}\left(\prod_{\ell=1}^{L}\lambda_{\ell}[i_{\ell}]\right)\left(y[i_{L}]\prod_{l=2}^{L}W_{\ell}[i_{\ell},i_{{\ell-1}}]\right)\cdot(W_{1}[i_{1}])^{\top}$ (9)
Let $\mathcal{P}_{k}=\{(p_{k},\ldots,p_{L}):p_{\ell}\in\{k,\ldots,d_{\ell}\}\}$ denote the set of all paths through layers $k,\ldots,L$, fix $p=(p_{1},\ldots,p_{L})$, i.e. $p\in\mathcal{P}_{1}$ and define the path activity function $\Lambda^{p}:\mathbb{R}^{d_{0}}\to[0,1]$ as:
$\Lambda^{p}$ $:=\prod_{\ell=1}^{L}\lambda_{\ell}[p_{\ell}]$ (10)
We may now rewrite the Equation 9 as:
$v_{\Lambda}(x)=\sum_{p\in\mathcal{P}_{1}}\Lambda^{p}(x)\cdot\left(y[p_{L}]\prod_{l=2}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}]\right)(W_{1}[p_{1}])^{\top}$ (12)
Note 3.1 Equation 12 expresses $v_{\Lambda}$ as a weighted sum of constant vector fields (we may call them atoms) indexed by $\mathcal{P}_{1}$ with the only dependence on $x$ being via $\Lambda^{p}$. In case of $\Lambda=G$ this amounts to filtering the atoms by a input-specific binary mask.
3.2 Tensor field pullbacks
Setting $P_{1}=|\mathcal{P}_{1}|=\prod_{\ell=1}^{L}d_{\ell}$ and some enumeration of $\mathcal{P}_{1}$ we may lift $\Lambda^{p}$ to the function $\tilde{\Lambda}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ by setting:
$\tilde{\Lambda}(x):=(\Lambda^{p}(x))_{p\in\mathcal{P}_{1}}$ (13)
By construction, $\tilde{\Lambda}(x)$ can be naturally identified with the tensor product of all $\lambda_{\ell}(x)$:
$\tilde{\Lambda}(x)\cong\bigotimes_{\ell=1}^{L}\lambda_{\ell}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}\cong\mathbb{R}^{P_{1}}$ (14)
hence we will call $\mathbb{R}^{P_{1}}$ a tensor product and refer to it’s elements as tensors. Similarly $\tilde{\Lambda}$ will be called a tensor field over $\mathbb{R}^{d_{0}}$.
More generally, for any piecewise continuous tensor field $\tau:\mathbb{R}^{d_{0}}\to[0,1]^{P_{1}}\subset\mathbb{R}^{P_{1}}$ we can define $\tau$-pullback $v_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{0}}$ as the following vector field:
$v_{\tau}(x):=\sum_{p\in\mathcal{P}_{1}}\tau_{p}(x)\cdot\left(y[p_{L}]\prod_{l=2}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}]\right)(W_{1}[p_{1}])^{\top}$ (15)
i.e. $v_{\tau}(x)$ is the sum of atoms with weights determined by the tensor $\tau(x)\in\mathbb{R}^{P_{1}}$. Consequently, we may define the $\tau$-induced function $f_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}$ as
$f_{\tau}(x):=\langle v_{\tau}(x),x\rangle$ (16)
It’s easy to see that $v_{\Lambda}=v_{\tilde{\Lambda}}$ and $f_{\Lambda}=f_{\tilde{\Lambda}}$ (by Equations 712 and 13), so we may safely overload the notation.
Note 3.2 $\tilde{\Lambda}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ is a rank-1 (i.e. pure) tensor, while $\tau(x)$ is allowed to be any tensor.
3.3 Path space and the linear nature of ReLU networks
Consider $\mathcal{P}_{0}$ i.e. the set of paths through layers $0,\ldots,L$ and $p=(p_{0},p_{1},\ldots,p_{L})$. Let $p|_{k}:=(p_{k},\ldots,p_{L})$, i.e. the (unique) sub-path of $p$ starting at layer $k$. We will abuse the notation and write $\tau_{p}$ instead of $\tau_{p|_{1}}$. With that remark we can expand $f_{\tau}(x)$ as:
$f_{\tau}(x)$ $=\langle v_{\tau},x\rangle=\sum_{p\in\mathcal{P}_{1}}\tau_{p}(x)\cdot\left(y[p_{L}]\prod_{l=2}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}]\right)\langle(W_{1}[p_{1}])^{\top},x\rangle$
$=\sum_{p\in\mathcal{P}_{1}}\tau_{p}(x)\cdot\left(y[p_{L}]\prod_{l=2}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}]\right)\left(\sum_{i=1}^{d_{0}}W_{1}[p_{1},i]\cdot x[i]\right)$
$=\sum_{p\in\mathcal{P}_{1}}\sum_{i=1}^{d_{0}}\tau_{p}(x)\cdot x[i]\cdot\left(y[p_{L}]\prod_{l=2}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}]\right)W_{1}[p_{1},i]$
$=\sum_{p\in\mathcal{P}_{0}}\tau_{p}(x)\cdot x[p_{0}]\cdot\left(y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}]\right)$
Again, we call $\mathbb{R}^{P_{0}}$ a tensor product as $\mathbb{R}^{P_{0}}\cong\bigotimes_{\ell=0}^{L}\mathbb{R}^{d_{\ell}}\cong(\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}})\otimes\mathbb{R}^{d_{0}}$. Additionally, we refer to $\mathbb{R}^{P_{0}}$ as the path space, as this is the space of all the neural paths in the network.
Let’s define $\Omega:\mathbb{R}^{d_{L}}\times\mathbb{R}^{d_{net}}\to\mathbb{R}^{P_{0}}$ as:
$\Omega((y,M)):=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (17)
and a $\tau$-induced feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ as:
$\phi_{\tau}(x):=(\tau_{p}\cdot x[p_{0}])_{p\in\mathcal{P}_{0}}\cong\tau\otimes x$ (18)
We may now express $f_{\tau}$ as a dot product in $\mathbb{R}^{P_{0}}$:
$f_{\tau}(x)=\sum_{p\in\mathcal{P}_{0}}(\phi_{\tau}(x))_{p}\cdot\omega_{p}=\langle\phi_{\tau}(x),\omega\rangle$ (19)
Observe that the dimension of $\mathbb{R}^{P_{0}}$ scales multiplicatively with the network’s depth and therefore, in deeper networks, it is orders of magnitude larger than the number of network parameters (which scales additively with depth). In a sense this means that neural networks, generally considered to be overparameterised models, are actually significantly underparameterised from the path space point of view.
Notice that the $\tau$-pullback defined earlier (Equation 15) is trivially equal to the pullback of $\omega^{\top}:\mathbb{R}^{P_{0}}\to\mathbb{R}$ through the feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$, because, by Equations 16 and 19, we have:
$\omega^{\top}(\phi_{\tau}(x))=\langle\omega,\phi_{\tau}(x)\rangle=f_{\tau}(x)=\langle v_{\tau}(x),x\rangle$ (20)
Note 3.3 The equation 20 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$, and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$. This disentangled view suggests that most of the network’s expressive power is due to the highly non-linear gating function $G$, as the network acts linearly (via $\omega\in\mathbb{R}^{P_{0}}$) on inputs transformed by $\phi_{\tilde{G}}$. Moreover, by Equation 20, the action of $\omega$ on $\phi_{\tilde{G}}(x)$ can be directly visualised in the input space via the action of the pullback $v_{G}$ on $x$.
Even though $\phi_{\tilde{G}^{(t)}}$ does change during training, it is significantly more stable than the direct mapping $M_{G}^{(t)}:x\mapsto x_{L}^{(t)}$, highlighting the importance of the path-representation.
4 Excitation pullback and the implicit linear model
Earlier work has demonstrated that only a small subnetwork within a larger model is responsible for most of its performance. In particular, weight pruning techniques demonstrate that many weights in trained networks can be removed without significantly degrading accuracy, implying the presence of sparse, efficient subnetworks within dense architectures (Han et al., 2015). Moreover, the Lottery Ticket Hypothesis (Frankle and Carbin, 2019) posits that such performant subnetworks often emerge early during training and can be “rewound” to their initial weights to train successfully in isolation. Together, these observations motivate the idea that the gating patterns in ReLU networks may approximately stabilize early in training, and thus may carry semantically meaningful structure even before convergence. Further motivation is provided in Section 6.1.
Let $X\subset\mathbb{R}^{d_{0}}$ be a fixed, compact dataset (e.g. $[0,1]^{d_{0}}$ in case of images) and let $g,h:\mathbb{R}^{d_{0}}\to\mathbb{R}^{n}$ be two functions. Define:
$\rho(x):=\operatorname{corr}(g(x),h(x))$
where $\operatorname{corr}$ denotes the Pearson correlation coefficient between the vectors $g(x)$ and $h(x)$. We say that $g$ and $h$ are positively correlated (in $X$) if $\rho(x)$ is positive for almost every $x\in X$, with respect to the Lebesgue measure. Similarly, we’ll call the correlation strong if $\rho(x)$ is high (e.g., greater than 0.7, although the exact threshold is not important) almost everywhere in $X$. Our core hypothesis is as follows:
Hypothesis 1 (Early emergence of stable excitation patterns).
There exists a binary tensor field $\gamma:\mathbb{R}^{d_{0}}\to\{0,1\}^{\mathcal{P}1}$ and an early training time $t\gamma\ll T$ such that for every $t>t_{\gamma}$, $\gamma$ is positively correlated with $\tilde{G}^{(t)}$ (the gating tensor field at time $t$).
Moreover, we assume that the network encodes the field $\gamma$ in its pre-activations. That is, there exists a gating function $\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}}\subset\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ such that:
  • $\Gamma$ is strongly positively correlated with the pre-activation function $Z:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$,
  • the induced rank-1 tensor field $\tilde{\Gamma}:\mathbb{R}^{d_{0}}\to[0,1]^{\mathcal{P}1}$ is strongly positively correlated with $\gamma$.
Comments and intuition:
  1. 1.
    We assume that the field $\gamma$ has binary weights and therefore acts as a path filter, just like $\tilde{G}$; however, it is not necessarily a rank-1 tensor, unlike $\tilde{G}$.
  2. 2.
    The field $\tilde{\Gamma}$ is composed of rank-1 tensors and thus can be seen as a uniform rank-1 approximation of $\gamma$.
  3. 3.
    Intuitively, $\gamma$ contains paths that pass through highly excited neurons (those with strong pre-activations).
  4. 4.
    The field $\tilde{\Gamma}$ provides a smoother approximation to $\gamma$ than any hard-gated field $\tilde{G}^{(t)}$, since it performs a form of soft thresholding; in particular, it suppresses barely active paths (i.e. those passing through weakly activated neurons) and partially recovers paths that are nearly active (i.e. those interrupted by a single weak neuron, but otherwise strong). This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and more aligned with the coarser structure of the decision boundary.
  5. 5.
    The function $\Gamma$ can be approximated as $\sigma\circ Z$, where $\sigma$ is a sigmoid-like nonlinearity applied element-wise, e.g. sigmoid, softsign (with properly adjusted range), normal cumulative distribution function (normal CDF), etc. In general, $\sigma$ should be neuron-specific or at least layer-specific, since it depends on the pre-activation distribution, which may be different for each neuron. In practice, thanks to the BatchNorm layers (Ioffe and Szegedy, 2015), a shared $\sigma$ across all neurons often yields good empirical results (even though BatchNorm still differentiates between the neuron distributions via it’s affine parameters).
  6. 6.
    $\tilde{\Gamma}$ approximates the direction of $\gamma$ (in the sense of cosine similarity) in $\mathbb{R}^{P_{1}}$, rather than its exact $L_{2}$-magnitude, since $\tilde{\Gamma}_{p}\ll 1$ for most paths $p$.
We refer to $\Gamma$ as the excitation function of the network and to $v_{\Gamma}$ as the excitation pullback.
Note 4.1 If the Hypothesis 1 holds, then, starting from time $t_{\gamma}$, the training dynamics of the network closely resembles the dynamics of a linear model $f_{\gamma}$ with a fixed feature map $\phi_{\gamma}$ (see Equation 18), i.e. a kernel machine operating under the feature noise introduced by $\phi_{\tilde{G}^{(t)}}$. In particular, the decision boundary of that implicit linear model $f_{\gamma}$ can be closely approximated by the decision boundary of $f_{\Gamma}$, the latter being directly visualizable by the excitation pullback $v_{\Gamma}$ (see Section 5).
Thus, the postulated $\gamma(x)$, potentially being a higher-rank tensor, would grasp the essence of what stays the same across different rank-1 neural responses $G^{(t)}(x)$ to the same input stimuli $x$ at time $t$ during network training.
5 Empirical validation
The goal of this section is to validate Hypothesis 1 empirically. In particular, if excitation pullbacks indeed approximate the decision boundary of a more regular linear model, then they should be better aligned with the data manifold, which should lead to better perceptual alignment. Moreover, we’d expect the excitation pullback to be highly robust to the choice of $\sigma$ across neurons, layers, and even architectures. This is because we’re interested in approximating the direction of $\gamma$ (not its exact value), and we assume $\gamma$ to be a binary tensor in $\mathbb{R}^{P_{1}}$, so it suffices to assign consistently higher scores to non-zero coordinates, implying robustness to hyperparameter choices.
To this end, we focus on popular and representative classes of ReLU-based architectures pretrained on ImageNet and available via the torchvision library, such that it is straightforward to replace most or all occurrences of ReLU and MaxPool2d (see Section 5.3) layers with our alternative variants, using a single recursive function over the child modules. Based on these criteria, we selected ResNet50, VGG11_BN, and DenseNet121.
Notably, determining the optimal choice of $\sigma$ for every neuron or layer in a given network is outside the scope of this work. Instead, we seek a single global set of hyperparameters that performs reasonably well across all layers of a given model - and ideally, across all selected architectures. Our goal is to evaluate whether the resulting excitation pullbacks indeed exhibit a substantially improved perceptual alignment compared to standard gradients.
5.1 Experimental setup
To generate visualizations, we use the val split of the publicly available Imagenette dataset (Howard, 2019), a subset of ImageNet containing 10 easily recognizable classes. To simplify presentation, we select every other class from the dataset, i.e.: tench, cassette player, church, garbage truck, and golf ball.
For each visualization, we construct a batch containing one random image from each of the five selected classes. We then compute excitation pullbacks for the classification neuron $y_{c}$ corresponding to each class, producing a $\times 5$ 5 grid: each row corresponds to an input image, and each column to the target class for which the pullback is computed. We also compute the pullback for a randomly selected ImageNet class (“ostrich”), shown in the last column.
We repeat the same setup, but instead of computing a single excitation pullback, we perform a rudimentary projected gradient ascent toward the logit (pre-activation) of each class, along the excitation pullback. Specifically, at each iteration we take a gradient step of $L_{2}$-norm 202020, whose direction is given by the excitation pullback. At each step we project the perturbed image on the ball of radius $\epsilon=100$ centered at the original image. We iterate this process 101010 times and plot both the final perturbed images and their differences from the originals. For contrast, we repeat the same procedure using vanilla input gradients in place of excitation pullbacks. Note that due to the choice of $\epsilon$, the perturbations are roughly fixed after 555 iterations (see Figure 1), the other 555 serving to further refine the perturbations and make them visually clearer.
5.2 Technical details
We can easily compute the excitation pullback $v_{\Gamma}$ by first computing the right pre-activations $z_{\ell}$ via $M_{G}$ (doing the ordinary forward pass) and then multiplying ReLU gradients by $\sigma(z_{\ell})$ in the backward pass, i.e. computing surrogate gradients for ReLU layers. This can be achieved in PyTorch (Paszke et al., 2019) by defining the appropriate torch.autograd.Function, see Listing 1. Alternatively, the more flexible Forward Gradient Injection can be used (Otte, 2024).
[…]
We download the pretrained weights via torchvision.models library. Specifically, we use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True.222see Section 6.3 for more context. Gradients are visualised by torchvision.utils.make_grid(scale_each=True). We perform the gradient ascent with images rescaled to the [1,1][-1,1][ - 1 , 1 ] range (and then appropriately normalized before feeding to the model). We seed the dataloader with torch.Generator().manual_seed(314).
5.3 Excitation pullbacks via MaxPool2d layer
Most convolutional architectures use spatial non-linearities in the form of MaxPool2d layers which can be viewed as a generalization of ReLU, i.e. $\text{ReLU}(z)=\max(0,z)$ and both operations propagate gradients only through the local maxima. This motivates the application of surrogate gradients to these layers as well. Analogously to the excitation pullback modification for ReLU, we leave the forward pass of max pooling unchanged, but replace its gradient with that of a softmax pooling operation, using the strike-through trick. This indeed smoothens the pullbacks even further. Implementation details can be found in the Listing 2; in particular, we set the temperature parameter to 0.30.30.3, the same as in surrogate gradient for ReLU.
5.4 Plotting excitation pullbacks
We present the plots generated according to the procedure described in Section 5.1, i.e. Figures 23 and 4. As shown in the experiments, a single choice of the function $\sigma$ performs reasonably well across all considered architectures, indicating a substantial robustness of excitation pullbacks to hyperparameter selection. As noted in Section 5.2, the visualizations remain qualitatively similar across a wide range of $\sigma$ choices. Moreover, excitation pullbacks tend to highlight similar features across architectures, which suggests that the models learn comparable feature representations. Finally, the structure of the excitation pullbacks intuitively reflects the internal organization of each network, reinforcing our hypothesis that they indeed faithfully capture the underlying decision process of the model.
Figure 2: Top row: vanilla gradients; bottom row: excitation pullbacks. From left to right: ResNet50, VGG11_BN, DenseNet121.
Figure 3: Left: Image perturbations after 10-step projected gradient ascent along excitation pullbacks toward each of the classes (columns). Right: Difference between the perturbed and original images. From top to bottom: ResNet50, VGG11_BN, DenseNet121. One can clearly distinguish label-specific features highlighted by the model on every image. Compare with the same plots for vanilla gradients, Figure 4
6 Discussion
6.1 Related work
Similar approach to ReLU networks from the path perspective has appeared in (Yadav et al., 2024) where authors study the so called Deep Linearly Gated Networks (DLGN) in which the gating function $G$ is computed by a parallel linear model $M$ with independent set of parameters $\theta$. The authors show that such models are considerably more interpretable than standard DNN’s and retain much of their performance. In their previous works (Lakshminarayanan and Singh, 2021) the authors define Neural Path Features (NPF) which are reminiscent of our gating functions (see Equation 3) and argue that “almost all the information learnt by a DNN with ReLU activations is stored in the gates”. They introduce Deep Gated Networks (DGN) and provide more evidence in their later work that gate adaptation is the key to generalisation (Lakshminarayanan and Singh, 2020), most notably, that “the winning lottery is in the gating pattern”, which reinforces our Hypothesis 1.
In (Linse et al., 2024) it is shown that setting the negative slope of LeakyReLU considerably higher in the backward pass than the forward pass leads to much better aligned global feature vizualisations via activation maximization, which can be directly explained by our results as LeakyReLU is equivalent to setting constant excitation function for each of the real half-lines (e.g. 1 for positive values and 0.3 for non-positive). In (Horuz et al., 2025) they generalize this idea to ReLU’s with different surrogate gradients, which is even closer to our approach, indicating that particular selection of surrogate ReLU gradient (e.g. B-SiLU) consistently improves generalization performance. Combined with our findings this leads to the open question, whether computing excitation pullbacks (instead of vanilla gradients) during training can improve generalization, and to what extent. Intuitively, one should not differentiate through the excitation gates explicitly as otherwise they may saturate and make the network overfit to the training data.
Another closely related work concerns local feature accentuations(Hamblin et al., 2024) where authors obtain high-quality local feature visualizations that are specific to the seeded image and the target feature. Furthermore, they argue that these accentuations are processed by the model along it’s natural circuit. Even though the produced images are of remarkably high quality, the proposed method uses strong image regularizations during the gradient ascent (in particular, images need to be transformed to the frequency domain) and requires as much as 100 gradient steps. Furthermore, the faithfulness of the produced explanations is established intuitively, without referencing the training dynamics.
It has been observed that networks trained to be robust to adversarial perturbations yield gradients that are often aligned with human perception, a phenomenon first observed in (Tsipras et al., 2019) and then referred to as perceptually-aligned gradients (PAG) in (Kaur et al., 2019). In (Srinivas et al., 2024) the authors attribute the PAG property to off-manifold robustness, which leads input gradients to lie approximately on the data manifold. This is resonates with our observation that excitation pullback $v_{\Gamma}$ provides a smoother approximation of the model’s decision boundary then the standard, hard-gated gradients (see Comments in Section 4) and therefore, by extension, may better align with the actual data manifold.
Our work suggests that even standard pretrained networks contain highly structured, semantically meaningful representations - but these get obfuscated by what we interpret as inherent gradient noise. This observation aligns with the broader perspective that explainability is not absent in modern networks, but merely hidden, and seemingly recoverable by principled methods such as ours.
6.2 On the possible biological plausibility of the implicit linear model
Intuitively, the relative stability of the gating mechanism throughout training is reminiscent of the approximately fixed synaptic connections observed in biological neural systems. In this analogy, the discrete gating pattern can be thought of as defining a persistent connectivity structure - akin to neural connections - while the excitation function describes the level of activity (excitation) of each neuron, resembling the frequency-based coding of information.
One may even venture a more speculative interpretation: the excitation pullbacks could be loosely associated with internal subjective representations, i.e. qualia - subjective, localized manifestations of experience that emerge from the brain activity taken as a whole - in a way similar to how excitation pullbacks emerge from the gating pattern in ReLU networks. This may open exciting direction in neuroscience and the philosophy of mind.
6.3 Limitations
Our experiments were run on pretrained torchvision models downloaded with the flag pretrained=True. This flag is now deprecated in the recent versions of the library, in particular it returns the older ResNet50_Weights.IMAGENET1K_V1. We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights, at least when using a single $\sigma$ for all the neurons, which, as we’ve indicated earlier, might by an oversimplified approach. We speculate that this is because of different distribution of pre-activations in newer versions of the pretrained models.
The formal theory developed in this paper concerns specifically the ReLU networks. Intuitively, similar arguments should hold for other activation functions, in particular the the ones similar to ReLU, e.g. SiLU and GELU. Initial experiments for those architectures show promise but require certain refinement of the computation of $\Gamma$ to account for the more complex gradients. However, it’s not straightforward to apply our results to attention-based models.
7 Acknowledgments
This research was supported by a private grant. The reader might find it interesting to review the public weekly log from the research.

### v3

@@
-Tapping into the Black Box: Uncovering Aligned Representations in Pretrained Neural Networks
-Abstract
-In this paper we argue that ReLU networks learn an implicit linear model we can actually tap into.
-We describe that alleged model formally and show that we can approximately pull its decision boundary back to the input space with certain simple modification to the backward pass.
-The resulting gradients (called excitation pullbacks) reveal high-resolution input- and target-specific features of remarkable perceptual alignment on a number of popular ImageNet-pretrained deep architectures.
-This strongly suggests that neural networks do, in fact, rely on learned interpretable patterns that can be recovered after training.
-Thus, our findings may have profound implications for knowledge discovery and the development of dependable artificial systems.
+Pulling Back the Curtain on ReLU Networks
+Pulling Back the Curtain on ReLU Networks
+Maciej Satkiewicz
+314 Foundation, Kraków
+maciej.satkiewicz@314.foundation
+Since any ReLU network is piecewise affine, its hidden units can be characterized by their pullbacks through the active subnetwork, i.e., by their gradients (up to bias terms).
+However, gradients of deeper neurons are notoriously misaligned, which obscures the network’s internal representations.
+We posit that models do align gradients with data, yet this is concealed by the intrinsic noise of the ReLU hard gating.
+We validate this intuition by applying soft gating in the backward pass only, reducing the local impact of weakly excited neurons.
+The resulting modified gradients, which we call excitation pullbacks, exhibit striking perceptual alignment on a number of ImageNet-pretrained architectures, while the rudimentary pixel-space gradient ascent quickly produces easily interpretable input- and target-specific features.
+Inspired by these findings, we formulate the path stability hypothesis, claiming that the binary activation patterns largely stabilize during training and get encoded in the pre-activation distribution of the final model.
+When true, excitation pullbacks become aligned with the gradients of a kernel machine that mainly determines the network’s decision.
+This provides a theoretical justification for the apparent faithfulness of the feature attributions based on these pullbacks, potentially even leading to mechanistic interpretability of deeper models.
+Incidentally, we give a possible explanation for the effectiveness of Batch Normalization and Deep Features, together with a novel perspective on the network’s internal memory and generalization properties.
+We release the code and an interactive app for easier exploration of the excitation pullbacks.
 Figure 1: A rudimentary 5-step pixel-space gradient ascent guided by excitation pullbacks for pretrained ResNet50.
@@
 See Section 6 for details.
+Best viewed digitally.
 1 Introduction
-This work investigates gradient-based explanations in ReLU networks, with a focus on convolutional architectures commonly used in computer vision.
-While our empirical results and theoretical developments concentrate on standard ReLU networks trained on image classification tasks, the core arguments are sufficiently general to apply - with the appropriate modifications - to other activation functions and modalities.
-We introduce a simple but effective method for interpreting neural predictions via excitation pullbacks, which isolate the robust informative signal in neural activations from the noise introduced by local data peculiarities.
-We motivate this idea extensively and demonstrate its empirical effectiveness on the following ImageNet-pretrained ReLU models: ResNet50 (He et al., 2015), VGG11_BN (Simonyan and Zisserman, 2015), and DenseNet121 (Huang et al., 2018).
-These specific architectures were chosen for reasons detailed in Section 5.
-We also give compelling theoretical arguments backing the claim that excitation pullbacks are indeed faithful local approximations of the network’s decision boundary.
-In particular, we hypothesize that they reveal a simpler linear model approximated by the trained network, encoded in the pre-activation patterns.
-The paper outline is as follows:
-  • Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks are described;
-  • Section 3 defines tensor field pullbacks and shows that neural networks are linear in their path space;
-  • Section 4 defines excitation pullbacks and postulates that they approximate the decision boundary of an implicit kernel machine that is learned in the network’s pre-activation patterns;
-  • Section 5 shows that excitation pullbacks of selected architectures pretrained on the ImageNet are perceptually aligned and can be used to produce faithful and meaningful local feature accentuations by just a few steps of gradient ascent along their directions;
-  • Section 6 describes the related works, indicates a possible biological feasibility of presented methods and discusses their limitations.
-This paper thus bridges practical gradient visualization with theoretical insights into neural function, suggesting a path toward more transparent and robust deep models.
+Understanding how neural networks organize their internal representations remains a longstanding and fundamental challenge in deep learning.
+In this paper we report the results of a less common line of research, where our primary goal was not incremental leaderboard improvement, but rather breaking through conceptual barriers.
+We’ve published111github.com/314-Foundation/ResearchLog the weekly research log to facilitate the discussion on this style of scientific inquiry.
+While we reduce the problem space to ReLU convolutional networks for vision tasks, our methods likely generalize to other activations with similar gating behavior, and potentially offer insights into a broader range of architectures and modalities.
+We originally set out by experimenting with prototypical-part based models such as ProtoPNet (Chen et al., 2019).
+However, this method of constructing self-explainable networks seems to have an unredeemable limitation in that it hinges on the black-box encoder network.
+To address this, we shifted our focus toward simpler, mechanistically interpretable architectures on CIFAR-10, with the goal of achieving gradient-data alignment.
+These experiments revealed an intriguing phenomenon: avoiding spatial gradient interference leads to substantially more interpretable gradients.
+Through a series of ablation studies, we then discovered that it is in fact possible to recover interpretable gradients in standard convolutional architectures by simply modifying the backward pass, leaving the forward computation intact.
+This led us to introduce what we call excitation pullbacks.
+The method turned out to scale to widely used ImageNet-pretrained architectures - including non-ReLU ones.
+However, we intentionally restrict the scope of this work to the following architectures: ResNet50 (He et al., 2015), VGG11_BN (Simonyan and Zisserman, 2015), and DenseNet121 (Huang et al., 2018); for reasons detailed in Section 6.
+The striking perceptual alignment of excitation pullbacks strongly suggests a deeper cause.
+Our primary concern was to validate theoretically that the produced explanations are reasonably faithful to the network’s decision process.
+Building on our previous investigations we developed a formal framework that allows us to analyze the pullbacks from the path perspective, where path is any sequence of neurons selected from the consecutive layers.
+In particular, we observe that ReLU networks correspond to linear models in the path space (the Hilbert space of formal linear combinations of all paths), under a feature map given by the tensor product of the binary activation vectors (encoding the on/off states of neurons) of subsequent layers (see Section 3 for details).
+This made us suspect that these feature maps might largely stabilize during training in the form of a higher-rank tensor field, which we call a path tensor feature (see Hypothesis 1).
+If that’s indeed the case, then the network behaves like a piecewise linear kernel machine with excitation pullbacks approximating directionally it’s gradient, so that the feature attributions they produce can be considered faithful to the network’s internal circuitry.
+This would translate to excitation pullbacks constituting a mechanistic interpretation of deep models in a style reminiscent of the “neural circuits” program (Cammarata et al., 2020) - but considerably streamlined by the pullback computation.
+Incidentally, the above considerations provide a possible explanation for the widespread effectiveness of Batch Normalization (Ioffe and Szegedy, 2015) in accelerating training, as normalizing neural activations significantly speeds up the stabilization of activation patterns and, consequently, the emergence of the path tensor feature and the global linear structure of the network’s internal representations, see Section 5.3.
+Furthermore, the success of Deep Features (Zhang et al., 2018) as a perceptual metric could be attributed to its kernel structure and the close relation to the path tensor feature (see Section 5.2).
+Interestingly, path tensor features allow us to view neural networks as associative memory systems, with excitation pullbacks as their contents (Section 5.4).
+Our work demonstrates that standard pretrained networks do in fact align gradients with data, and highlights how one can extract interpretable representations of hidden neurons.
+Furthermore, the proposed formal framework sheds light on the internal organization of neural networks, and even offers a potential explanation for their generalization capabilities: if the stability condition expressed in Hypothesis 1 is satisfied, then the network’s decisions can be well approximated by a kernel machine whose gradient is feasibly accessible by the excitation pullback.
+The striking perceptual alignment observed in our experiments strongly suggests that this may indeed be the case, at least to some extent.
+The paper is organized as follows:
+Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks; Section 3 defines tensor field pullbacks and shows that ReLU networks correspond to linear maps in the path space; Section 4 introduces the excitation pullbacks and discusses the inherent gradient noise; Section 5 discusses the path stability hypothesis and its implications; Section 6 details the pullback visualization method used in the paper; Section 7 describes the related works and discusses the limitations of presented approach.
+We release the code222github.com/314-Foundation/ExcitationPullbacks and the interactive demo app333 huggingface.co/spaces/msat/ExcitationPullbacks to make the exploration of excitation pullbacks easily accessible.
 2 Preliminaries and notational conventions
@@
 2.1 ReLU networks
-Let $M$ be a feedforward ReLU network with parameters $\theta\in\mathbb{R}^{d_{net}}$, expressible as a composition of affine transformations and pointwise ReLU nonlinearities.111we omit max-pooling layers from the analysis to improve clarity, while noting that similar arguments are likely to apply due to the structural similarity between max-pooling and ReLU; see Section 5.3.
+Let $M$ be a feedforward ReLU network with parameters $\theta\in\mathbb{R}^{d_{net}}$, expressible as a composition of affine transformations and pointwise ReLU nonlinearities.444we omit max-pooling layers from the analysis to improve clarity, noting that similar arguments are likely to apply due to the structural similarity between max-pooling and ReLU; see Section 6.3.
 Let $L$ denote the number of layers, and let $d_{\ell}$ be the width of the $\ell$-th layer, with $d_{0}$ denoting the input dimension.
@@
 The network applies the following operations:
-$z_{\ell}=W_{\ell}x_{\ell-1}+b_{\ell},\quad x_{\ell}=\text{ReLU}(z_{\ell}),\qquad\text{for }\ell=1,\dotsc,L$
+$z_{\ell}:=W_{\ell}x_{\ell-1}+b_{\ell},\quad x_{\ell}:=\text{ReLU}(z_{\ell}),\qquad\text{for }\ell=1,\dotsc,L$
 where $W_{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$, $b_{\ell}\in\mathbb{R}^{d_{\ell}}$, and $\text{ReLU}(z)=\max(0,z)$ is applied element-wise.
-We define the overall network function $M:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ by the recursive application of the layers, i.e.,
-$M(x_{0}):=x_{L}$
+We identify $M$ with the overall network function $x_{L}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$, i.e., $M(x):=x_{L}(x)$.
+We’ll often abuse the notation and write $x_{\ell}\in\mathbb{R}^{d_{\ell}}$, meaning the value of $x_{\ell}(x)$ for some fixed $x\in\mathbb{R}^{d_{0}}$; same for $z_{\ell}$ and other similar functions.
 In this paper we’ll treat $M$ as backbone for some linear model, i.e.
@@
 Standard classification network with $C$ classes has $C$ corresponding vectors $y_{c}$.
-We will skip the lower index $c$ assuming that it’s fixed.
+We will usually skip the lower index $c$ assuming that it’s fixed.
 For notational brevity, we consider augmented matrices and vectors where the bias term is incorporated by appending a 1 to the end of the input vector.
 Specifically, we define the augmented input and weight matrix as
-$\tilde{x}_{\ell-1}=\begin{bmatrix}x_{\ell-1}\\ 1\end{bmatrix}\in\mathbb{R}^{d_{\ell-1}+1},\quad\tilde{W}_{\ell}=\begin{bmatrix}W_{\ell}&b_{\ell}\\ 0\quad\cdots\quad 0&1\end{bmatrix}\in\mathbb{R}^{(d_{\ell}+1)\times(d_{\ell-1}+1)}$
+$\tilde{x}_{\ell-1}:=\begin{bmatrix}x_{\ell-1}\\ 1\end{bmatrix}\in\mathbb{R}^{d_{\ell-1}+1},\quad\tilde{W}_{\ell}:=\begin{bmatrix}W_{\ell}&b_{\ell}\\ 0\quad\cdots\quad 0&1\end{bmatrix}\in\mathbb{R}^{(d_{\ell}+1)\times(d_{\ell-1}+1)}$
 Then, the augmented pre-activation vector is given by
-$\tilde{z}_{\ell}=\tilde{W}_{\ell}\tilde{x}_{\ell-1}$
+$\tilde{z}_{\ell}:=\tilde{W}_{\ell}\tilde{x}_{\ell-1}$
 Conventions
@@
 We will occasionally write $\mathbb{R}^{d^{*}_{0}}\subset\mathbb{R}^{d_{0}}$ to denote the true input space (without the implicitly added coordinate).
-Notice that, by construction, the $\mathbb{R}^{d^{*}_{0}}$ remains the actual domain of $f$.
+Notice that $\mathbb{R}^{d^{*}_{0}}$ remains the actual domain of $f$, as the extra coordinate is fixed.
 2.2 Gating representation of ReLU networks
@@
 $x_{\ell}=g_{\ell}\odot z_{\ell}$
-where $g_{\ell}\in\{0,1\}^{d_{\ell}}$ is a binary gating vector defined by
+where $\odot$ is the Hadamard product and $g_{\ell}\in\{0,1\}^{d_{\ell}}$ is a binary gating vector defined by
 $g_{\ell}[i]=\begin{cases}1&\text{if }z_{\ell}[i]>0,\\ 0&\text{otherwise}.\end{cases}$
-Since $z_{\ell}$ is itself a function of the input $x_{0}$, we may define a gating function
-$G:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\{0,1\}^{d_{\ell}},\quad G(x):=(g_{1},\dots,g_{L})$ (1)
-Let $G_{\ell}(x)=\mathrm{diag}(g_{\ell})\in\mathbb{R}^{d_{\ell}\times d_{\ell}}$ denote the diagonal matrix whose diagonal entries are given by $g_{\ell}$.
-Then, the network output $M(x)=x_{L}$ can be written as a product of (augmented) weight matrices interleaved with input-dependent diagonal matrices:
+Since $g_{\ell}$ is itself a function of the input $x$, we may define the gating function:
+$G:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\{0,1\}^{d_{\ell}},\quad G(x):=(g_{1}(x),\dots,g_{L}(x))$ (1)
+Let $G_{\ell}(x)=\mathrm{diag}(g_{\ell}(x))\in\{0,1\}^{d_{\ell}\times d_{\ell}}$ denote the diagonal matrix whose diagonal entries are given by $g_{\ell}(x)$.
+Then, the network output can be written as a product of (augmented) weight matrices interleaved with input-dependent diagonal matrices:
 $M(x)=G_{L}(x)W_{L}G_{L-1}(x)W_{L-1}\cdots G_{1}(x)W_{1}x$
 where each $W_{\ell}$ includes the bias terms via the convention introduced earlier.
-This representation makes the piecewise-linear structure of ReLU networks explicit: for a fixed input $x$, the gating pattern $(g_{1},\dots,g_{L})\in\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ determines a purely linear computation.
+This representation makes the piecewise-linear structure of ReLU networks explicit: for a fixed input $x$, the gating pattern $(g_{1},\dots,g_{L})\in\prod_{\ell=1}^{L}\{0,1\}^{d_{\ell}}$ determines a purely linear computation.
 Similarly, let’s define the pre-activation function as:
-$Z:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}},\quad Z(x):=(z_{1},\dots,z_{L})$ (2)
+$Z:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}},\quad Z(x):=(z_{1}(x),\dots,z_{L}(x))$ (2)
 i.e. the concatenation of all pre-activations of $M(x)$.
@@
 $M_{\Lambda}(x):=\Lambda_{L}(x)W_{L}\Lambda_{L-1}(x)W_{L-1}\cdots\Lambda_{1}(x)W_{1}x$ (4)
-where $\Lambda_{\ell}(x)=\mathrm{diag}(\lambda_{\ell}(x))\in\mathbb{R}^{d_{\ell}\times d_{\ell}}$ is the diagonal matrix formed from the $\ell$-th component $\lambda_{\ell}$ of $\Lambda(x)$.
+where $\Lambda_{\ell}(x)=\mathrm{diag}(\lambda_{\ell}(x))\in[0,1]^{d_{\ell}\times d_{\ell}}$ is the diagonal matrix formed from $\lambda_{\ell}(x)$.
 In particular, $M$ is $G$-induced, i.e.
@@
 We can pull the linear map $y^{\top}:\mathbb{R}^{d_{L}}\to\mathbb{R}$ back to the input space $\mathbb{R}^{d_{0}}$ through $M_{\Lambda}$, generating the following covector field over $\mathbb{R}^{d_{0}}$:
-$v^{\top}_{\Lambda}:=y^{\top}\vec{M}_{\Lambda}$
-We will call both $v^{\top}_{\Lambda}$ and the corresponding vector field $v_{\Lambda}$ a $\Lambda$-pullback. In particular, we have:
-$v_{G}^{\top}=y^{\top}\vec{M}_{G}$ (6)
-Now, the following sequence of equations can be easily verified:
+$v^{\top}_{\Lambda}:=y^{\top}\vec{M}_{\Lambda}$ (6)
+We will call both $v^{\top}_{\Lambda}$ and the corresponding vector field $v_{\Lambda}$ a $\Lambda$-pullback.
+In particular, we have $v_{G}^{\top}=y^{\top}\vec{M}_{G}$.
+The following sequence of equations can be easily verified:
 $f_{\Lambda}(x)=y^{\top}M_{\Lambda}(x)=y^{\top}\vec{M}_{\Lambda}(x)x=v_{\Lambda}^{\top}(x)x=\langle v_{\Lambda}(x),x\rangle$ (7)
@@
 2.6 Training dynamics
-[…]
-3 Neural networks are linear in their path space
+Network parameters $\theta$ and $u$ change during training.
+We index time steps by $t\in\{0,\ldots,T\}$, where $t=0$ denotes the initialization and $t=T$ corresponds to the final state of training.
+Whenever an upper index $(t)$ appears, it indicates the value of the corresponding object at training step $t$, for example $(y^{(t)},\theta^{(t)})$ denotes the network parameters and $G^{(t)}$ denotes the gating function induced by $M^{(t)}$.
+For notational brevity, we omit the index $t=T$ when referring to the final trained model.
+3 The path space and tensor field pullbacks
 3.1 Network action as atom filtering
@@
 the one corresponding to the $i$-th output in the $\ell$-th layer of $M$.
-Due to the compositional nature of the network, we can define the pullback $v_{\Lambda}|_{(\ell,i)}$ of $u_{\ell}[i]$ in the same way we did for $y$ just consider the appropriate subnetwork below $u_{\ell}[i]$ as the backbone.
+Due to the compositional nature of the network, we can define the pullback $v_{\Lambda}|_{(\ell,i)}$ of $u_{\ell}[i]$ in the same way we did for $y$ - just consider the appropriate subnetwork below $u_{\ell}[i]$ as the backbone.
 Moreover, the full-network pullback $v_{\Lambda}$ can be expressed as a weighted sum of the pullbacks of neurons from previous layers, for example:
 $v_{\Lambda}=\sum_{i=1}^{d_{L}}y[i]\cdot\lambda_{L}[i]\cdot v_{\Lambda}|_{(L,i)}$
-[…]
+i.e.
+$v_{\Lambda}$ is a weighted sum of layer-$L$ pullbacks.
+This means that neurons in layer $L$ contribute particular directions $v_{\Lambda}|_{(L,i)}$ to the full pullback $v_{\Lambda}$, weighted by $\lambda_{L}[i_{L}]$, giving an explicit sense to the intuition that neurons form a compositional hierarchy of feature detectors.
 We can also express $v_{\Lambda}$ as the weighted sum of first-layer pullbacks:
 $v_{\Lambda}=\sum_{(i_{1},i_{2},\ldots,i_{L})}y[i_{L}]\cdot\lambda_{L}[i_{L}]\cdot W_{L}[i_{L},i_{L-1}]\cdots\lambda_{2}[i_{2}]\cdot W_{2}[i_{2},i_{1}]\cdot\lambda_{1}[i_{1}]\cdot v_{\Lambda}|_{(1,i_{1})}$
-Where $(i_{1},\ldots,i_{L})$ is the path through layers $1,\ldots,L$, i.e. the selection of indices $i_{\ell}\in\{1,\ldots,d_{\ell}\}$. Now rearrange the factors:
+Where $(i_{1},\ldots,i_{L})$ is the path through layers $1,\ldots,L$, i.e. the selection of indices $i_{\ell}\in\{1,\ldots,d_{\ell}\}$. Now let’s rearrange the factors:
 $v_{\Lambda}=\sum_{(i_{1},i_{2},\ldots,i_{L})}\left(\prod_{\ell=1}^{L}\lambda_{\ell}[i_{\ell}]\right)\left(y[i_{L}]\prod_{l=2}^{L}W_{\ell}[i_{\ell},i_{{\ell-1}}]\right)\cdot v_{\Lambda}|_{(1,i_{1})}$
@@
 $v_{\Lambda}=\sum_{(i_{1},i_{2},\ldots,i_{L})}\left(\prod_{\ell=1}^{L}\lambda_{\ell}[i_{\ell}]\right)\left(y[i_{L}]\prod_{l=2}^{L}W_{\ell}[i_{\ell},i_{{\ell-1}}]\right)\cdot(W_{1}[i_{1}])^{\top}$ (9)
-Let $\mathcal{P}_{k}=\{(p_{k},\ldots,p_{L}):p_{\ell}\in\{k,\ldots,d_{\ell}\}\}$ denote the set of all paths through layers $k,\ldots,L$, fix $p=(p_{1},\ldots,p_{L})$, i.e.
-$p\in\mathcal{P}_{1}$ and define the path activity function $\Lambda^{p}:\mathbb{R}^{d_{0}}\to[0,1]$ as:
-$\Lambda^{p}$ $:=\prod_{\ell=1}^{L}\lambda_{\ell}[p_{\ell}]$ (10)
+Let $\mathcal{P}_{k}=\{(p_{k},\ldots,p_{L}):p_{\ell}\in\{k,\ldots,d_{\ell}\}\}$ denote the set of all paths through layers $k,\ldots,L$, fix $p=(p_{1},\ldots,p_{L})\in\mathcal{P}_{1}$ and define the path activity function $\Lambda^{p}:\mathbb{R}^{d_{0}}\to[0,1]$ as:
+$\Lambda^{p}(x):=\prod_{\ell=1}^{L}\lambda_{\ell}(x)[p_{\ell}]$ (10)
 We may now rewrite the Equation 9 as:
 $v_{\Lambda}(x)=\sum_{p\in\mathcal{P}_{1}}\Lambda^{p}(x)\cdot\left(y[p_{L}]\prod_{l=2}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}]\right)(W_{1}[p_{1}])^{\top}$ (11)
-Note 3.1 Equation 12 expresses $v_{\Lambda}$ as a weighted sum of constant vector fields (we may call them atoms) indexed by $\mathcal{P}_{1}$ with the only dependence on $x$ being via $\Lambda^{p}$.
-In case of $\Lambda=G$ this amounts to filtering the atoms by a input-specific binary mask.
+Note 3.1 Equation 11 expresses $v_{\Lambda}$ as a weighted sum of constant vector fields (atoms) indexed by $\mathcal{P}_{1}$, with the only dependence on $x$ being via $\Lambda^{p}$.
+In case of $\Lambda=G$ this amounts to filtering the atoms by an input-specific binary mask.
 3.2 Tensor field pullbacks
-Setting $P_{1}=|\mathcal{P}_{1}|=\prod_{\ell=1}^{L}d_{\ell}$ and some enumeration of $\mathcal{P}_{1}$ we may lift $\Lambda^{p}$ to the function $\tilde{\Lambda}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ by setting:
+Setting $P_{1}=|\mathcal{P}_{1}|=\prod_{\ell=1}^{L}d_{\ell}$ and choosing some enumeration of $\mathcal{P}_{1}$ we may lift $\Lambda^{p}$ to the function $\tilde{\Lambda}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ as:
 $\tilde{\Lambda}(x):=(\Lambda^{p}(x))_{p\in\mathcal{P}_{1}}$ (12)
@@
 $\tilde{\Lambda}(x)\cong\bigotimes_{\ell=1}^{L}\lambda_{\ell}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}\cong\mathbb{R}^{P_{1}}$ (13)
-hence we will call $\mathbb{R}^{P_{1}}$ a tensor product and refer to it’s elements as tensors. Similarly $\tilde{\Lambda}$ will be called a tensor field over $\mathbb{R}^{d_{0}}$.
+We will call $\mathbb{R}^{P_{1}}$ a tensor product and refer to it’s elements as tensors. Similarly, $\tilde{\Lambda}$ will be called a tensor field over $\mathbb{R}^{d_{0}}$.
 More generally, for any piecewise continuous tensor field $\tau:\mathbb{R}^{d_{0}}\to[0,1]^{P_{1}}\subset\mathbb{R}^{P_{1}}$ we can define $\tau$-pullback $v_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{0}}$ as the following vector field:
@@
 $f_{\tau}(x):=\langle v_{\tau}(x),x\rangle$ (15)
-It’s easy to see that $v_{\Lambda}=v_{\tilde{\Lambda}}$ and $f_{\Lambda}=f_{\tilde{\Lambda}}$ (by Equations 712 and 13), so we may safely overload the notation.
-Note 3.2 $\tilde{\Lambda}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ is a rank-1 (i.e. pure) tensor, while $\tau(x)$ is allowed to be any tensor.
+It’s easy to see that $v_{\Lambda}=v_{\tilde{\Lambda}}$ and $f_{\Lambda}=f_{\tilde{\Lambda}}$ (Equations 711 and 12), so we may safely overload the notation.
+Note 3.2 $\tilde{\Lambda}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ is a rank-1 tensor, while $\tau(x)$ can have arbitrary rank.
 3.3 Path space and the linear nature of ReLU networks
-Consider $\mathcal{P}_{0}$ i.e.
-the set of paths through layers $0,\ldots,L$ and $p=(p_{0},p_{1},\ldots,p_{L})$.
-Let $p|_{k}:=(p_{k},\ldots,p_{L})$, i.e.
-the (unique) sub-path of $p$ starting at layer $k$.
+Consider $\mathcal{P}_{0}$, the set of all paths through layers $0,\ldots,L$ and $p=(p_{0},p_{1},\ldots,p_{L})\in\mathcal{P}_{0}$.
+Let $p|_{k}:=(p_{k},\ldots,p_{L})$ denote the (unique) sub-path of $p$ starting at layer $k$.
 We will abuse the notation and write $\tau_{p}$ instead of $\tau_{p|_{1}}$.
@@
 Again, we call $\mathbb{R}^{P_{0}}$ a tensor product as $\mathbb{R}^{P_{0}}\cong\bigotimes_{\ell=0}^{L}\mathbb{R}^{d_{\ell}}\cong(\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}})\otimes\mathbb{R}^{d_{0}}$.
-Additionally, we refer to $\mathbb{R}^{P_{0}}$ as the path space, as this is the space of all the neural paths in the network.
+Additionally, we refer to $\mathbb{R}^{P_{0}}$ as the path space, as this is the space of all neural paths in the network.
 Let’s define $\Omega:\mathbb{R}^{d_{L}}\times\mathbb{R}^{d_{net}}\to\mathbb{R}^{P_{0}}$ as:
-$\Omega((y,M)):=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (17)
+$\Omega((y,\theta)):=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (16)
 and a $\tau$-induced feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ as:
@@
 In a sense this means that neural networks, generally considered to be overparameterised models, are actually significantly underparameterised from the path space point of view.
-Notice that the $\tau$-pullback defined earlier (Equation 15) is trivially equal to the pullback of $\omega^{\top}:\mathbb{R}^{P_{0}}\to\mathbb{R}$ through the feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$, because, by Equations 16 and 19, we have:
+Notice that the $\tau$-pullback defined earlier (Equation 14) is trivially equal to the pullback of $\omega^{\top}:\mathbb{R}^{P_{0}}\to\mathbb{R}$ through the feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$, because, by Equations 15 and 18, we have:
 $\omega^{\top}(\phi_{\tau}(x))=\langle\omega,\phi_{\tau}(x)\rangle=f_{\tau}(x)=\langle v_{\tau}(x),x\rangle$ (19)
-Note 3.3 The equation 20 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$, and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$.
+Note 3.3 The equation 19 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$ and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$.
 This disentangled view suggests that most of the network’s expressive power is due to the highly non-linear gating function $G$, as the network acts linearly (via $\omega\in\mathbb{R}^{P_{0}}$) on inputs transformed by $\phi_{\tilde{G}}$.
 Moreover, by Equation 19, the action of $\omega$ on $\phi_{\tilde{G}}(x)$ can be directly visualised in the input space via the action of the pullback $v_{G}$ on $x$.
-Even though $\phi_{\tilde{G}^{(t)}}$ does change during training, it is significantly more stable than the direct mapping $M_{G}^{(t)}:x\mapsto x_{L}^{(t)}$, highlighting the importance of the path-representation.
-4 Excitation pullback and the implicit linear model
+Even though $\phi_{\tilde{G}^{(t)}}$ does change during training, it is significantly more stable than the direct mapping $M_{G}^{(t)}:x\mapsto x_{L}^{(t)}\in\mathbb{R}^{d_{L}}$, hinting at the importance of this representation.
+4 Excitation pullbacks and the inherent gradient noise
+4.1 On the inherent gradient noise
+Despite the following representation property (see Equations 7 and 8):
+$f(x)$ $=\langle v_{G}(x),x\rangle=v_{G}^{\top}(x)x$
+$\nabla f(x)$ $=v_{G}^{\top}(x)|_{\mathbb{R}^{d_{0}^{*}}}$
+the gradients of ReLU networks are notoriously hard to interpret (see Figure 2).
+This may suggest that those models rely on some strange predictive patterns that are meaningless to humans.
+But it may also mean that gradients are contaminated by an inherent noise that obfuscates a more regular underlying decision boundary.
+In this section, we advocate the latter view.
+Indeed, by Equation 6 the pullback vector $v_{G}(x)$, and hence also the gradient $\nabla f(x)$, depend solely on the network parameters $(y,\theta)$ and the gating pattern $G(x)$.
+This means that $v_{G}(x)$ looses a lot of information about $x$, in particular about the pre-activation values $z_{\ell}(x)$.
+This loss of information might be important for the training process, as the network is less likely to overfit, but it is problematic for interpretability during inference.
+To highlight the potential problems, suppose that $x\in\mathbb{R}^{d_{0}}$ lies on the boundary of two linear regions of $M$.
+Let $x_{a}\sim x$ and $x_{b}\sim x$ be two nearby points lying on the opposite sides of the boundary, i.e.
+each in a different linear region.
+By continuity of $f$, we have
+$f(x_{a})\approx f(x)\approx f(x_{b})$
+which, by Equation 7, translates to
+$\langle v_{G}(x_{a}),x_{a}\rangle\approx\langle v_{G}(x),x\rangle\approx\langle v_{G}(x_{b}),x_{b}\rangle.$
+By properties of the dot product we get
+$\langle v_{G}(x_{a}),x\rangle\approx\langle v_{G}(x_{b}),x\rangle\implies\langle v_{G}(x_{a})-v_{G}(x_{b}),x\rangle\approx 0$ (20)
+This means that, in general, the pullback vectors $v_{G}(x_{a})$ and $v_{G}(x_{b})$ may differ significantly as long as their difference remains orthogonal to $x$, i.e.
+it can be any point from the the entire hyperplane orthogonal to $x$.
+In standard training there is little incentive to align these orthogonal directions as they hardly affect the network’s value $f(x)$ (incidentally, such incentive may be provided by adversarial training, see Section 7.1).
+Note 4.1 Intuitively, $v_{G}(x)$ may contain locally irrelevant directions from neurons that are spuriously active, i.e.
+barely above the activation threshold but not corresponding to any salient feature in $x$.
+On the other hand, some important neurons may be inactive, either due to the local noise or because $x$ is a negative example and does not contain many features relevant to the chosen class $c$ for which the gradient is computed.
+4.2 Introducing excitation pullbacks
+The considerations from previous section suggest that we should look beyond the hard activation gates $G(x)$ to properly pull the action of $y^{\top}$ back to $\mathbb{R}^{d_{0}}$ - in a way that is robust to hyper-local noise.
+A natural idea is to perform soft gating, i.e.
+compose the pre-activations $z_{\ell}$ with some sigmoid-like step function $\sigma_{\ell}:\mathbb{R}^{d_{\ell}}\to[0,1]^{d_{\ell}}$:
+$\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}},\quad\Gamma:=(\gamma_{1},\dots,\gamma_{L}),\quad\gamma_{\ell}:=\sigma_{\ell}\circ z_{\ell}:\mathbb{R}^{d_{0}}\to[0,1]^{d_{\ell}}$ (21)
+In general, $\sigma_{\ell}$ is neuron-specific, since the pre-activation distribution may be different for each neuron.
+In practice, thanks to the BatchNorm layers, a shared $\sigma$ across all neurons yields good empirical results (even though BatchNorm still differentiates between the neuron distributions via it’s affine parameters).
+In the paper we set the global element-wise $\sigma_{\ell}=\sigma=\operatorname{sigmoid(\frac{z}{0.3})}:\mathbb{R}\to[0,1]$, but other options (e.g.
+rescaled softsign or Normal CFD) lead to very similar results - as long as the general sigmoid shape of the global $\sigma$ is preserved, see Section 6.2.
+Because $\sigma_{\ell}$ can be understood as measuring the excitation of a neuron we may accordingly call $\Gamma$ the excitation function, it’s corresponding pullback $v_{\Gamma}$ the excitation pullback, and it’s induced tensor field $\tilde{\Gamma}$ the excitation tensor field.
+Note that in order to compute $\Gamma$ we still need to use the original gates $G$ in the forward pass to compute the right pre-activations.
+Note 4.2 The tensor field $\tilde{\Gamma}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ (see Equation 12) suppresses barely active paths, i.e.
+those passing through weakly activated neurons; and partially recovers paths that are nearly active (i.e.
+those interrupted by a single weak neuron, but otherwise strong).
+This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and probably more aligned with the coarser structure of the decision boundary.
+In a sense, $v_{\Gamma}$ is dominated by atoms corresponding to highly excited paths, as it quickly suppresses the less-excited paths (see Equation 10).
+5 On the path stability and its potential significance
+5.1 The path stability hypothesis
+Because of the relative stability of $\phi_{\tilde{G}^{(t)}}$ (Note 3.3), it’s tempting to assume that the gating function $G^{(t)}$ gets fixed early on during training.
+We consider this assumption too optimistic as the gates probably do change simply because of the gradient noise we’ve hinted at before.
+Notice, however, that $\tilde{G}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ maps inputs to binary rank-1 tensors, meaning that it has limited path-filtering capability (Note 3.1).
+But perhaps there are more refined atom filters (corresponding to higher-rank binary tensors) that do get fixed early on during training?
+We motivate this assumption below.
 Earlier work has demonstrated that only a small subnetwork within a larger model is responsible for most of its performance.
@@
 Moreover, the Lottery Ticket Hypothesis (Frankle and Carbin, 2019) posits that such performant subnetworks often emerge early during training and can be “rewound” to their initial weights to train successfully in isolation.
-Together, these observations motivate the idea that the gating patterns in ReLU networks may approximately stabilize early in training, and thus may carry semantically meaningful structure even before convergence.
-Further motivation is provided in Section 6.1.
-Let $X\subset\mathbb{R}^{d_{0}}$ be a fixed, compact dataset (e.g. $[0,1]^{d_{0}}$ in case of images) and let $g,h:\mathbb{R}^{d_{0}}\to\mathbb{R}^{n}$ be two functions. Define:
-$\rho(x):=\operatorname{corr}(g(x),h(x))$
-where $\operatorname{corr}$ denotes the Pearson correlation coefficient between the vectors $g(x)$ and $h(x)$.
-We say that $g$ and $h$ are positively correlated (in $X$) if $\rho(x)$ is positive for almost every $x\in X$, with respect to the Lebesgue measure.
-Similarly, we’ll call the correlation strong if $\rho(x)$ is high (e.g., greater than 0.7, although the exact threshold is not important) almost everywhere in $X$.
-Our core hypothesis is as follows:
-Hypothesis 1 (Early emergence of stable excitation patterns).
-There exists a binary tensor field $\gamma:\mathbb{R}^{d_{0}}\to\{0,1\}^{\mathcal{P}1}$ and an early training time $t\gamma\ll T$ such that for every $t>t_{\gamma}$, $\gamma$ is positively correlated with $\tilde{G}^{(t)}$ (the gating tensor field at time $t$).
-Moreover, we assume that the network encodes the field $\gamma$ in its pre-activations.
-That is, there exists a gating function $\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}}\subset\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ such that:
-  • $\Gamma$ is strongly positively correlated with the pre-activation function $Z:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$,
-  • the induced rank-1 tensor field $\tilde{\Gamma}:\mathbb{R}^{d_{0}}\to[0,1]^{\mathcal{P}1}$ is strongly positively correlated with $\gamma$.
-Comments and intuition:
-  1. 1.
-    We assume that the field $\gamma$ has binary weights and therefore acts as a path filter, just like $\tilde{G}$; however, it is not necessarily a rank-1 tensor, unlike $\tilde{G}$.
-  2. 2.
-    The field $\tilde{\Gamma}$ is composed of rank-1 tensors and thus can be seen as a uniform rank-1 approximation of $\gamma$.
-  3. 3.
-    Intuitively, $\gamma$ contains paths that pass through highly excited neurons (those with strong pre-activations).
-  4. 4.
-    The field $\tilde{\Gamma}$ provides a smoother approximation to $\gamma$ than any hard-gated field $\tilde{G}^{(t)}$, since it performs a form of soft thresholding; in particular, it suppresses barely active paths (i.e.
-those passing through weakly activated neurons) and partially recovers paths that are nearly active (i.e.
-those interrupted by a single weak neuron, but otherwise strong).
-This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and more aligned with the coarser structure of the decision boundary.
-  5. 5.
-    The function $\Gamma$ can be approximated as $\sigma\circ Z$, where $\sigma$ is a sigmoid-like nonlinearity applied element-wise, e.g.
-sigmoid, softsign (with properly adjusted range), normal cumulative distribution function (normal CDF), etc.
-In general, $\sigma$ should be neuron-specific or at least layer-specific, since it depends on the pre-activation distribution, which may be different for each neuron.
-In practice, thanks to the BatchNorm layers (Ioffe and Szegedy, 2015), a shared $\sigma$ across all neurons often yields good empirical results (even though BatchNorm still differentiates between the neuron distributions via it’s affine parameters).
-  6. 6.
-    $\tilde{\Gamma}$ approximates the direction of $\gamma$ (in the sense of cosine similarity) in $\mathbb{R}^{P_{1}}$, rather than its exact $L_{2}$-magnitude, since $\tilde{\Gamma}_{p}\ll 1$ for most paths $p$.
-We refer to $\Gamma$ as the excitation function of the network and to $v_{\Gamma}$ as the excitation pullback.
-Note 4.1 If the Hypothesis 1 holds, then, starting from time $t_{\gamma}$, the training dynamics of the network closely resembles the dynamics of a linear model $f_{\gamma}$ with a fixed feature map $\phi_{\gamma}$ (see Equation 18), i.e.
-a kernel machine operating under the feature noise introduced by $\phi_{\tilde{G}^{(t)}}$.
-In particular, the decision boundary of that implicit linear model $f_{\gamma}$ can be closely approximated by the decision boundary of $f_{\Gamma}$, the latter being directly visualizable by the excitation pullback $v_{\Gamma}$ (see Section 5).
-Thus, the postulated $\gamma(x)$, potentially being a higher-rank tensor, would grasp the essence of what stays the same across different rank-1 neural responses $G^{(t)}(x)$ to the same input stimuli $x$ at time $t$ during network training.
+Furthermore, in (Lakshminarayanan and Singh, 2021) the authors define Neural Path Features (NPF), which are essentially our gating functions (Equation 3), and argue that “almost all the information learnt by a DNN with ReLU activations is stored in the gates”.
+Specifically, they introduce Deep Gated Networks (DGN) and provide empirical evidence that gate adaptation is the key to generalization, most notably, that “the winning lottery is in the gating pattern”  (Lakshminarayanan and Singh, 2020).
+Together, these observations motivate the idea that the gating patterns in ReLU networks may approximately stabilize early in training, and thus may carry semantically meaningful structure long before convergence.
+Recall that $f^{(t)}(x)=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle$ and $f^{(t)}_{\tau}(x)=\langle\omega^{(t)},\phi_{\tau}(x)\rangle$.
+Let $X=(x_{i})_{i}\subset\mathbb{R}^{d_{0}}$ be the training dataset and $f(X)=(f(x_{i}))_{i}$ be the vector of all the values of $f$.
+We say that two vectors $a,b\in\mathbb{R}^{n}$ are strongly positively correlated if the Pearson correlation coefficient $\rho(a,b)$ is high (e.g., greater than 0.9, although the exact threshold value is not important).
+We denote this relation as $a\approx_{+}b$.
+We now state our core hypothesis:
+Hypothesis 1 (Early emergence of stable paths).
+There exists a piecewise constant binary tensor field $\pi:\mathbb{R}^{d_{0}}\to\{0,1\}^{{P_{1}}}$ and an early training time $t_{\pi}\ll T$ such that for every $t>t_{\pi}$ we have $f^{(t)}(X)\approx_{+}f^{(t)}_{\pi}(X)$.
+Additionally, we postulate that the network encodes the tensor field $\pi$ in its pre-activation function $Z$.
+That is, there exists a choice of $(\sigma_{\ell})_{\ell}$ (as in Equation 21) such that for every $x\in X$ we have $\tilde{\Gamma}(x)\approx_{+}\pi(x)$.
+We refer to the above assumption as the path stability condition and to any such $\pi$ as path tensor feature.
+Intuition:
+The hypothesis states that, even though the gating tensors $\tilde{G}^{(t)}(x)$ change during training, there exists a time-independent gating tensor $\pi(x)$ that, starting from time $t_{\pi}$, largely determines network’s decision $f^{(t)}(x)$ at $x$.
+In this view, the variability of tensors $\tilde{G}^{(t)}(x)$ can be attributed to their limited expressive power as rank-1 tensors; the network, trying to approximate $\pi(x)$, learns it implicitly in it’s pre-activation pattern $Z(x)$ in the form of the set of highly-excited paths.
+The latter can be feasibly directionally approximated by the rank-1 non-binary tensor $\tilde{\Gamma}(x)$.
+5.2 The implicit kernel machine and the faithfulness of excitation pullbacks
+Recall that $\omega^{(t)}=\Omega((y^{(t)},\theta^{(t)}))\in\mathbb{R}^{P_{0}}$ (Equation 16) encodes the network’s parameters at time $t$ and that:
+$f^{(t)}(x)$ $=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle$
+$f^{(t)}_{\pi}(x)$ $=\langle\omega^{(t)},\phi_{\pi}(x)\rangle$
+If the Hypothesis 1 holds, then, starting from time $t_{\pi}$ the training dynamics of $\omega^{(t)}$ closely resembles the training dynamics of the kernel machine given by $f^{(t)}_{\pi}$.
+In fact, $(\phi_{\tilde{G}^{(t)}}-\phi_{\pi})$ can be interpreted as a feature noise added to the fixed feature map $\phi_{\pi}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ at time $t$ of training.
+Note 5.1 Hypothesis 1 implies that gradient descent actually learns the implicit kernel machine $f_{\pi}$ and the model $M$ is just a noisy wrapper that allows us to feasibly approximate the extremely-high-dimensional feature function $\phi_{\pi}$.
+Since $\pi$ is assumed to be piecewise constant, then also it’s pullback $v_{\pi}$ is piecewise constant and, by Equation 15, we have
+$\nabla f_{\pi}(x)=v_{\pi}^{\top}(x)|_{\mathbb{R}^{d_{0}^{*}}}$ (22)
+This means that the pullback $v_{\pi}^{\top}$ can actually be interpreted as the gradient of the kernel machine $f_{\pi}$.
+We can feasibly approximate the direction of that gradient by $v_{\Gamma}$, which is a better approximation than that of $v_{G}$ and therefore $v_{\Gamma}^{\top}$ is actually more faithful to the network’s decision than the standard gradient (if the Hypothesis 1 holds).
+Also notice that in order to compute the path space dot product $\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle$ it suffices to compute the following product kernel:
+$K_{\Lambda}(x,x^{\prime}):=\langle x,x^{\prime}\rangle\prod_{\ell=1}^{L}\langle\lambda_{\ell}(x),\lambda_{\ell}(x^{\prime})\rangle=\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle,\quad\text{for }x,x^{\prime}\in\mathbb{R}^{d_{0}}$ (23)
+In particular, we may feasibly study the properties of $\langle\phi_{\pi}(x),\phi_{\pi}(x^{\prime})\rangle$ with the help of $K_{\Gamma}$ kernel.
+Notice that $K_{\Gamma}$ encodes the idea behind Deep Features and thus offers a possible explanation for why the latter is an effective perceptual metric (Zhang et al., 2018), as it basically expresses the path-space dot product under the feature map $\phi_{\pi}$ (up to the scaling), tapping into the internal representation of $f_{\pi}$.
+5.3 The effectiveness of Batch Normalization
+Batch Normalization (Ioffe and Szegedy, 2015) has become one of the most influential architectural innovations in modern deep learning, largely due to its ability to dramatically accelerate training and improve generalization.
+While the standard explanation attributes these benefits to reduced internal covariate shift and smoother optimization landscapes (Santurkar et al., 2019), our framework offers an alternative and complementary perspective.
+Note that BN basically removes the bias from inputs and largely prevents the emergence of data-dominant output neurons, enforcing their uniform activation across the batch.
+This facilitates both the neuron specialization and more uniform usage of the layer width.
+Highly activated neurons are therefore considerably better aligned (directionally) with the particular input then the less activated ones and have larger impact on the weight update during training, which reinforces their role.
+All these phenomena favor quicker stabilization of path activations, i.e.
+faster emergence of the path tensor feature and the corresponding linear structure of network representations, improving the training speed and representational quality.
+Interestingly, this resonates with the recent line of work on B-cos networks (Böhle et al., 2022), which explicitly promote weight-input alignment as a mechanism for interpretability.
+The alignment of excitation pullbacks suggests that the role of B-cos transform is already performed by the BN layer in standard architectures, but this is obfuscated by the pullbacks of weakly excited neurons contaminating the gradient, see Section 4.1.
+5.4 ReLU network as content-addressable storage
+Under the path stability hypothesis, ReLU networks can be interpreted as an associative memory system with excitation pullbacks as its content.
+Recall the Note 3.1 where we interpret the pullback vector field $v_{G}$ as binary filtering of atoms.
+This can be understood as computing the target location $\tilde{G}(x)\in\mathbb{R}^{\mathcal{P}_{1}}$ in the network’s memory space $\mathbb{R}^{\mathcal{P}_{1}}$.
+During training, for a positive example $x$, the network updates its weights to pull $v_{G}^{(t)}(x)$ closer to $x$ (due to Equation 7), i.e.
+it combines $x$ with the value $v_{G^{(t)}}(x)$ stored at the location $\tilde{G}^{(t)}(x)$.
+Under the path stability condition this location remains largely the same, oscillating around $\pi(x)$ and therefore $v_{\pi}(x)$ (and $v_{\Gamma}(x)$) contains the most coherent information.
+This is validated by the perceptual alignment of excitation pullbacks.
+During inference, the network first computes the location $\tilde{G}(x)$ and then takes the dot product of $x$ with the stored memories $v_{(c,G)}(x)$ for every class $c$, returning the class of the best-matching memory.
+According to the path stability hypothesis, we can extract the actual memories by computing $v_{(c,\Gamma)}(x)$ instead (the excitation pullback for class $c$).
+It may however change the network’s prediction if the pullbacks weren’t used during the training.
 6 Empirical validation
-The goal of this section is to validate Hypothesis 1 empirically.
-In particular, if excitation pullbacks indeed approximate the decision boundary of a more regular linear model, then they should be better aligned with the data manifold, which should lead to better perceptual alignment.
-Moreover, we’d expect the excitation pullback to be highly robust to the choice of $\sigma$ across neurons, layers, and even architectures.
-This is because we’re interested in approximating the direction of $\gamma$ (not its exact value), and we assume $\gamma$ to be a binary tensor in $\mathbb{R}^{P_{1}}$, so it suffices to assign consistently higher scores to non-zero coordinates, implying robustness to hyperparameter choices.
+This section describes how we visualize the excitation pullbacks.
 To this end, we focus on popular and representative classes of ReLU-based architectures pretrained on ImageNet and available via the torchvision library, such that it is straightforward to replace most or all occurrences of ReLU and MaxPool2d (see Section 6.3) layers with our alternative variants, using a single recursive function over the child modules.
 Based on these criteria, we selected ResNet50, VGG11_BN, and DenseNet121.
-Notably, determining the optimal choice of $\sigma$ for every neuron or layer in a given network is outside the scope of this work.
+Notably, determining the optimal choice of $\sigma_{\ell}$ for every neuron or layer in a given network is outside the scope of this work.
 Instead, we seek a single global set of hyperparameters that performs reasonably well across all layers of a given model - and ideally, across all selected architectures.
@@
 For each visualization, we construct a batch containing one random image from each of the five selected classes.
-We then compute excitation pullbacks for the classification neuron $y_{c}$ corresponding to each class, producing a $\times 5$ 5 grid: each row corresponds to an input image, and each column to the target class for which the pullback is computed.
+We then compute excitation pullbacks for the classification neuron $y_{c}$ corresponding to each class, producing a $5\times 5$ grid: each row corresponds to an input image, and each column to the target class for which the pullback is computed.
 We also compute the pullback for a randomly selected ImageNet class (“ostrich”), shown in the last column.
 We repeat the same setup, but instead of computing a single excitation pullback, we perform a rudimentary projected gradient ascent toward the logit (pre-activation) of each class, along the excitation pullback.
-Specifically, at each iteration we take a gradient step of $L_{2}$-norm 202020, whose direction is given by the excitation pullback.
+We perform the gradient ascent with images rescaled to the [1,1][-1,1] range (and then appropriately normalized before feeding to the model).
+Specifically, at each iteration we take a gradient step of $L_{2}$-norm $20$, whose direction is given by the excitation pullback.
 At each step we project the perturbed image on the ball of radius $\epsilon=100$ centered at the original image.
-We iterate this process 101010 times and plot both the final perturbed images and their differences from the originals.
+We iterate this process $10$ times and plot both the final perturbed images and their differences from the originals.
 For contrast, we repeat the same procedure using vanilla input gradients in place of excitation pullbacks.
-Note that due to the choice of $\epsilon$, the perturbations are roughly fixed after 555 iterations (see Figure 1), the other 555 serving to further refine the perturbations and make them visually clearer.
+Note that due to the choice of $\epsilon$, the perturbations are roughly fixed after $5$ iterations (see Figure 1), the other $5$ serving to further refine the perturbations and make them visually clearer.
 6.2 Technical details
-We can easily compute the excitation pullback $v_{\Gamma}$ by first computing the right pre-activations $z_{\ell}$ via $M_{G}$ (doing the ordinary forward pass) and then multiplying ReLU gradients by $\sigma(z_{\ell})$ in the backward pass, i.e.
+We can easily compute the excitation pullback $v_{\Gamma}$ by first computing the right pre-activations $z_{\ell}$ via $M_{G}$ (doing the ordinary forward pass) and then multiplying ReLU gradients by $\sigma_{\ell}(z_{\ell})$ in the backward pass, i.e.
 computing surrogate gradients for ReLU layers.
@@
 Alternatively, the more flexible Forward Gradient Injection can be used (Otte, 2024).
-[…]
-We download the pretrained weights via torchvision.models library.
-Specifically, we use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True.222see Section 6.3 for more context.
+In the experiments we use a fixed global $\sigma$ for all the layers in all the networks, but ideally $\sigma_{\ell}$ could be adjusted for every hidden neuron separately, based on the neuron pre-activation distribution.
+We’ve observed that different choices of the global $\sigma$ produced similarly-looking pullback visualizations as long as the overall sigmoidal shape of the global $\sigma$ function remained the same.
+We’ve chosen $\sigma(z)=\operatorname{sigmoid(\frac{z}{temp})}$ for $\operatorname{temp}=0.3$ but any selection of the parameter temp from the approximate range $\text{temp}\in[0.15,0.5]$ seemed to work quite well across all the tested architectures, which implies the considerable robustness of the method to the hyperparameter choice.
+Notice that $\Gamma$ approaches hard gating $G$ as temp goes to zero and no gating at all as temp goes to infinity.
+We use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True from the torchvision.models library.555see Section 7.3 for more context.
 Gradients are visualised by torchvision.utils.make_grid(scale_each=True).
-We perform the gradient ascent with images rescaled to the [1,1][-1,1][ - 1 , 1 ] range (and then appropriately normalized before feeding to the model).
 We seed the dataloader with torch.Generator().manual_seed(314).
-5.3 Excitation pullbacks via MaxPool2d layer
+6.3 Excitation pullbacks through the MaxPool2d layer
 Most convolutional architectures use spatial non-linearities in the form of MaxPool2d layers which can be viewed as a generalization of ReLU, i.e.
@@
 This indeed smoothens the pullbacks even further.
-Implementation details can be found in the Listing 2; in particular, we set the temperature parameter to 0.30.30.3, the same as in surrogate gradient for ReLU.
+Implementation details can be found in the Listing 2; we set the temperature parameter to $0.3$, the same as in surrogate gradient for ReLU.
 6.4 Plotting excitation pullbacks
@@
 Figures 23 and 4.
-As shown in the experiments, a single choice of the function $\sigma$ performs reasonably well across all considered architectures, indicating a substantial robustness of excitation pullbacks to hyperparameter selection.
+As shown in the experiments, a single global choice of the function $\sigma$ performs reasonably well across all considered architectures, indicating a substantial robustness of excitation pullbacks to hyperparameter selection.
 As noted in Section 6.2, the visualizations remain qualitatively similar across a wide range of $\sigma$ choices.
@@
 Finally, the structure of the excitation pullbacks intuitively reflects the internal organization of each network, reinforcing our hypothesis that they indeed faithfully capture the underlying decision process of the model.
-Figure 2: Top row: vanilla gradients; bottom row: excitation pullbacks. From left to right: ResNet50, VGG11_BN, DenseNet121.
-Figure 3: Left: Image perturbations after 10-step projected gradient ascent along excitation pullbacks toward each of the classes (columns).
-Right: Difference between the perturbed and original images.
-From top to bottom: ResNet50, VGG11_BN, DenseNet121.
-One can clearly distinguish label-specific features highlighted by the model on every image.
-Compare with the same plots for vanilla gradients, Figure 4
+Figure 2: Top row: vanilla gradients; bottom row: excitation pullbacks. From left to right: ResNet50, VGG11_BN, DenseNet121. Best viewed digitally.
 7 Discussion
 7.1 Related work
-Similar approach to ReLU networks from the path perspective has appeared in (Yadav et al., 2024) where authors study the so called Deep Linearly Gated Networks (DLGN) in which the gating function $G$ is computed by a parallel linear model $M$ with independent set of parameters $\theta$.
-The authors show that such models are considerably more interpretable than standard DNN’s and retain much of their performance.
-In their previous works (Lakshminarayanan and Singh, 2021) the authors define Neural Path Features (NPF) which are reminiscent of our gating functions (see Equation 3) and argue that “almost all the information learnt by a DNN with ReLU activations is stored in the gates”.
-They introduce Deep Gated Networks (DGN) and provide more evidence in their later work that gate adaptation is the key to generalisation (Lakshminarayanan and Singh, 2020), most notably, that “the winning lottery is in the gating pattern”, which reinforces our Hypothesis 1.
+Similar approach to ReLU networks from the path perspective has appeared in (Yadav et al., 2024) where authors study the so-called Deep Linearly Gated Networks (DLGN) in which the gating function $G$ is computed by a parallel linear model $M$ with independent set of parameters $\theta$.
+They show that such models are considerably more interpretable than standard DNN’s and retain much of their performance.
+We highlight the authors’ other relevant contributions in Section 5.1, motivating our Hypothesis 1.
+Overall, they arrive at similar formal constructions throughout their works but their main focus is on alternative architectures like DGNs; in contrast, we present a refined formal framework to study properties of standard architectures.
 In (Linse et al., 2024) it is shown that setting the negative slope of LeakyReLU considerably higher in the backward pass than the forward pass leads to much better aligned global feature vizualisations via activation maximization, which can be directly explained by our results as LeakyReLU is equivalent to setting constant excitation function for each of the real half-lines (e.g.
@@
 Combined with our findings this leads to the open question, whether computing excitation pullbacks (instead of vanilla gradients) during training can improve generalization, and to what extent.
-Intuitively, one should not differentiate through the excitation gates explicitly as otherwise they may saturate and make the network overfit to the training data.
+Intuitively, one should not differentiate through the excitation gates explicitly as they may saturate and overfit; nevertheless, they can still be computed in the backward pass as alternative gradients.
 Another closely related work concerns local feature accentuations(Hamblin et al., 2024) where authors obtain high-quality local feature visualizations that are specific to the seeded image and the target feature.
@@
 In (Srinivas et al., 2024) the authors attribute the PAG property to off-manifold robustness, which leads input gradients to lie approximately on the data manifold.
-This is resonates with our observation that excitation pullback $v_{\Gamma}$ provides a smoother approximation of the model’s decision boundary then the standard, hard-gated gradients (see Comments in Section 4) and therefore, by extension, may better align with the actual data manifold.
-Our work suggests that even standard pretrained networks contain highly structured, semantically meaningful representations - but these get obfuscated by what we interpret as inherent gradient noise.
-This observation aligns with the broader perspective that explainability is not absent in modern networks, but merely hidden, and seemingly recoverable by principled methods such as ours.
+This resonates with our observation that excitation pullback $v_{\Gamma}$ provides a smoother approximation of the model’s decision boundary then the standard, hard-gated gradients and therefore, by extension, may better align with the actual data manifold.
 7.2 On the possible biological plausibility of the implicit linear model
@@
 One may even venture a more speculative interpretation: the excitation pullbacks could be loosely associated with internal subjective representations, i.e.
-qualia - subjective, localized manifestations of experience that emerge from the brain activity taken as a whole - in a way similar to how excitation pullbacks emerge from the gating pattern in ReLU networks.
+qualia - subjective, localized manifestations of experience that emerge from the brain activity taken as a whole - in a way similar to how excitation pullbacks emerge from the gating patterns in ReLU networks.
+It makes even more sense in the light of Section 5.4, where we interpret excitation pullbacks as the content of the network’s associative memory.
 This may open exciting direction in neuroscience and the philosophy of mind.
@@
 This flag is now deprecated in the recent versions of the library, in particular it returns the older ResNet50_Weights.IMAGENET1K_V1.
-We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights, at least when using a single $\sigma$ for all the neurons, which, as we’ve indicated earlier, might by an oversimplified approach.
-We speculate that this is because of different distribution of pre-activations in newer versions of the pretrained models.
+We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights, at least when using a single global $\sigma$ for all the neurons, which, as we’ve indicated earlier, might by an oversimplified approach.
+We speculate that this is because of different distribution of pre-activations in newer versions of the pretrained models, which have been heavily regularized.
 The formal theory developed in this paper concerns specifically the ReLU networks.
-Intuitively, similar arguments should hold for other activation functions, in particular the the ones similar to ReLU, e.g.
+Intuitively, similar arguments should hold for other activation functions, in particular the ones similar to ReLU, e.g.
 SiLU and GELU.
 Initial experiments for those architectures show promise but require certain refinement of the computation of $\Gamma$ to account for the more complex gradients.
-However, it’s not straightforward to apply our results to attention-based models.
+We haven’t yet explored the potential applicability of our method to attention-based models.
+Additionally, we expect the convolutions to play rather significant role in the emergence of aligned excitation pullbacks, as spatial weight sharing seems to be crucial for the early path stabilization.
+This underlines the importance of an appropriate inductive bias for our results to carry over to other domains.
 8 Acknowledgments
-This research was supported by a private grant. The reader might find it interesting to review the public weekly log from the research.
+This research was conducted at the 314 Foundation [^6] https://314.foundation with support from a private grant awarded specifically for this study.

### v4

@@
 When true, excitation pullbacks become aligned with the gradients of a kernel machine that mainly determines the network’s decision.
-This provides a theoretical justification for the apparent faithfulness of the feature attributions based on these pullbacks, potentially even leading to mechanistic interpretability of deeper models.
+This provides a theoretical justification for the apparent faithfulness of the feature attributions based on excitation pullbacks, potentially even leading to mechanistic interpretability of deep models.
 Incidentally, we give a possible explanation for the effectiveness of Batch Normalization and Deep Features, together with a novel perspective on the network’s internal memory and generalization properties.
@@
 In this paper we report the results of a less common line of research, where our primary goal was not incremental leaderboard improvement, but rather breaking through conceptual barriers.
-We’ve published111github.com/314-Foundation/ResearchLog the weekly research log to facilitate the discussion on this style of scientific inquiry.
+We’ve published [^1] https://github.com/314-Foundation/ResearchLog the weekly research log to facilitate the discussion on this style of scientific inquiry.
 While we reduce the problem space to ReLU convolutional networks for vision tasks, our methods likely generalize to other activations with similar gating behavior, and potentially offer insights into a broader range of architectures and modalities.
@@
 The striking perceptual alignment of excitation pullbacks strongly suggests a deeper cause.
-Our primary concern was to validate theoretically that the produced explanations are reasonably faithful to the network’s decision process.
-Building on our previous investigations we developed a formal framework that allows us to analyze the pullbacks from the path perspective, where path is any sequence of neurons selected from the consecutive layers.
-In particular, we observe that ReLU networks correspond to linear models in the path space (the Hilbert space of formal linear combinations of all paths), under a feature map given by the tensor product of the binary activation vectors (encoding the on/off states of neurons) of subsequent layers (see Section 3 for details).
+Our primary concern was to motivate theoretically that the produced explanations are reasonably faithful to the network’s decision process, as this is the main unsolved issue in the attribution research (Deng et al., 2025).
+Building on our investigations we developed a formal framework that allows us to analyze the pullbacks from the path perspective, where path is any sequence of neurons selected from the consecutive layers.
+In particular, we’ve observed that ReLU networks correspond to linear models in the path space (the Hilbert space of formal linear combinations of all paths), under a feature map given by the tensor product of the binary activation vectors (encoding the on/off states of neurons) of subsequent layers (see Section 3 for details).
 This made us suspect that these feature maps might largely stabilize during training in the form of a higher-rank tensor field, which we call a path tensor feature (see Hypothesis 1).
 If that’s indeed the case, then the network behaves like a piecewise linear kernel machine with excitation pullbacks approximating directionally it’s gradient, so that the feature attributions they produce can be considered faithful to the network’s internal circuitry.
-This would translate to excitation pullbacks constituting a mechanistic interpretation of deep models in a style reminiscent of the “neural circuits” program (Cammarata et al., 2020) - but considerably streamlined by the pullback computation.
-Incidentally, the above considerations provide a possible explanation for the widespread effectiveness of Batch Normalization (Ioffe and Szegedy, 2015) in accelerating training, as normalizing neural activations significantly speeds up the stabilization of activation patterns and, consequently, the emergence of the path tensor feature and the global linear structure of the network’s internal representations, see Section 5.3.
-Furthermore, the success of Deep Features (Zhang et al., 2018) as a perceptual metric could be attributed to its kernel structure and the close relation to the path tensor feature (see Section 5.2).
-Interestingly, path tensor features allow us to view neural networks as associative memory systems, with excitation pullbacks as their contents (Section 5.4).
-Our work demonstrates that standard pretrained networks do in fact align gradients with data, and highlights how one can extract interpretable representations of hidden neurons.
-Furthermore, the proposed formal framework sheds light on the internal organization of neural networks, and even offers a potential explanation for their generalization capabilities: if the stability condition expressed in Hypothesis 1 is satisfied, then the network’s decisions can be well approximated by a kernel machine whose gradient is feasibly accessible by the excitation pullback.
+This would translate to excitation pullbacks constituting a mechanistic interpretation of deep models in a style reminiscent of the “neural circuits” program (Cammarata et al., 2020), but considerably streamlined by the pullback computation.
+We view our hypothesis primarily as a lens through which neural networks can be studied, rather than a statement that is universally valid.
+We expect it to hold only partially, but nonetheless consider its formulation useful for guiding further research.
+In particular, it inspires a possible explanation for the widespread effectiveness of Batch Normalization (Section 5.3) and the success of Deep Features as a perceptual metric (Section 5.2).
+Even more interestingly, it allows us to view classification networks as associative memory systems, with excitation pullbacks as their contents (Section 5.4).
+Our work demonstrates that standard pretrained networks do seem to align gradients with data, and highlights how one can extract interpretable and arguably faithful representations of hidden neurons.
+Furthermore, the proposed formal framework sheds light on the internal organization of neural networks, and even offers a potential explanation for their generalization capabilities: if the path stability condition expressed in Hypothesis 1 is satisfied, then the network’s decisions can be well approximated by a kernel machine whose gradient is feasibly computable by the excitation pullback.
 The striking perceptual alignment observed in our experiments strongly suggests that this may indeed be the case, at least to some extent.
+While our formal contributions remain preliminary and partly speculative, we see them as pointing toward a potentially rich avenue for future work.
 The paper is organized as follows:
-Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks; Section 3 defines tensor field pullbacks and shows that ReLU networks correspond to linear maps in the path space; Section 4 introduces the excitation pullbacks and discusses the inherent gradient noise; Section 5 discusses the path stability hypothesis and its implications; Section 6 details the pullback visualization method used in the paper; Section 7 describes the related works and discusses the limitations of presented approach.
-We release the code222github.com/314-Foundation/ExcitationPullbacks and the interactive demo app333 huggingface.co/spaces/msat/ExcitationPullbacks to make the exploration of excitation pullbacks easily accessible.
+Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks; Section 3 defines tensor field pullbacks and shows that ReLU networks correspond to linear maps in the path space; Section 4 introduces the excitation pullbacks and discusses the inherent gradient noise; Section 5 presents the path stability hypothesis and its implications; Section 6 details the pullback visualization method used in the paper; Section 7 describes the related works and discusses the limitations of presented approach.
+We release the code [^2] https://github.com/314-Foundation/ExcitationPullbacks and the interactive demo app [^3] https://huggingface.co/spaces/msat/ExcitationPullbacks to make the exploration of excitation pullbacks easily accessible.
 2 Preliminaries and notational conventions
@@
 $\tilde{z}_{\ell}:=\tilde{W}_{\ell}\tilde{x}_{\ell-1}$
-Conventions
+Conventions.
 Throughout the paper, we will often abuse notation by omitting the tilde and the explicit increase of dimension, writing simply
@@
 This construction allows us to view $f$ as a scalar function defined directly on the domain $\mathbb{R}^{d_{0}}$ (through the action of $\phi$).
-Convention
+Conventions.
 Whenever a piecewise continuous vector field $v:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{0}}$ satisfies
@@
 By construction, $\tilde{\Lambda}(x)$ can be naturally identified with the tensor product of all $\lambda_{\ell}(x)$:
-$\tilde{\Lambda}(x)\cong\bigotimes_{\ell=1}^{L}\lambda_{\ell}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}\cong\mathbb{R}^{P_{1}}$ (13)
+$\tilde{\Lambda}(x)=\bigotimes_{\ell=1}^{L}\lambda_{\ell}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}\cong\mathbb{R}^{P_{1}}$ (13)
 We will call $\mathbb{R}^{P_{1}}$ a tensor product and refer to it’s elements as tensors. Similarly, $\tilde{\Lambda}$ will be called a tensor field over $\mathbb{R}^{d_{0}}$.
@@
 Let’s define $\Omega:\mathbb{R}^{d_{L}}\times\mathbb{R}^{d_{net}}\to\mathbb{R}^{P_{0}}$ as:
-$\Omega((y,\theta)):=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (16)
+$\Omega((y,\theta)):=\omega=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (16)
 and a $\tau$-induced feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ as:
-$\phi_{\tau}(x):=(\tau_{p}\cdot x[p_{0}])_{p\in\mathcal{P}_{0}}\cong\tau\otimes x$ (17)
+$\phi_{\tau}(x):=(\tau_{p}(x)\cdot x[p_{0}])_{p\in\mathcal{P}_{0}}=\tau(x)\otimes x$ (17)
 We may now express $f_{\tau}$ as a dot product in $\mathbb{R}^{P_{0}}$:
-$f_{\tau}(x)=\sum_{p\in\mathcal{P}_{0}}(\phi_{\tau}(x))_{p}\cdot\omega_{p}=\langle\phi_{\tau}(x),\omega\rangle$ (18)
+$f_{\tau}(x)=\sum_{p\in\mathcal{P}_{0}}(\phi_{\tau}(x))_{p}\cdot\omega_{p}=\langle\phi_{\tau}(x),\omega\rangle=\langle\tau(x)\otimes x,\omega\rangle$ (18)
 Observe that the dimension of $\mathbb{R}^{P_{0}}$ scales multiplicatively with the network’s depth and therefore, in deeper networks, it is orders of magnitude larger than the number of network parameters (which scales additively with depth).
-In a sense this means that neural networks, generally considered to be overparameterised models, are actually significantly underparameterised from the path space point of view.
+In a sense, this let’s us consider neural networks as underparameterised models on the path space.
 Notice that the $\tau$-pullback defined earlier (Equation 14) is trivially equal to the pullback of $\omega^{\top}:\mathbb{R}^{P_{0}}\to\mathbb{R}$ through the feature function $\phi_{\tau}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$, because, by Equations 15 and 18, we have:
 $\omega^{\top}(\phi_{\tau}(x))=\langle\omega,\phi_{\tau}(x)\rangle=f_{\tau}(x)=\langle v_{\tau}(x),x\rangle$ (19)
-Note 3.3 The equation 19 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$ and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$.
+Note 3.3 Equation 18 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$, and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$.
 This disentangled view suggests that most of the network’s expressive power is due to the highly non-linear gating function $G$, as the network acts linearly (via $\omega\in\mathbb{R}^{P_{0}}$) on inputs transformed by $\phi_{\tilde{G}}$.
@@
 A natural idea is to perform soft gating, i.e.
-compose the pre-activations $z_{\ell}$ with some sigmoid-like step function $\sigma_{\ell}:\mathbb{R}^{d_{\ell}}\to[0,1]^{d_{\ell}}$:
-$\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}},\quad\Gamma:=(\gamma_{1},\dots,\gamma_{L}),\quad\gamma_{\ell}:=\sigma_{\ell}\circ z_{\ell}:\mathbb{R}^{d_{0}}\to[0,1]^{d_{\ell}}$ (21)
-In general, $\sigma_{\ell}$ is neuron-specific, since the pre-activation distribution may be different for each neuron.
+compose the pre-activations $z_{\ell}$ with some sigmoid-like step functions $\sigma_{\ell,i}:\mathbb{R}\to[0,1]$ applied element-wise:
+$\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}},\quad\Gamma:=(\gamma_{1},\dots,\gamma_{L}),\quad\gamma_{\ell}:=\sigma_{\ell}\circ z_{\ell}:\mathbb{R}^{d_{0}}\to[0,1]^{d_{\ell}},\quad\sigma_{\ell}(z_{\ell}):=\bigl(\sigma_{\ell,i}(z_{\ell}[i])\bigr)_{i=1}^{d_{\ell}}$ (21)
+In general, we use neuron-specific functions $\sigma_{\ell,i}$, since the pre-activation distribution may be different for each neuron.
 In practice, thanks to the BatchNorm layers, a shared $\sigma$ across all neurons yields good empirical results (even though BatchNorm still differentiates between the neuron distributions via it’s affine parameters).
-In the paper we set the global element-wise $\sigma_{\ell}=\sigma=\operatorname{sigmoid(\frac{z}{0.3})}:\mathbb{R}\to[0,1]$, but other options (e.g.
+In the paper we use a global element-wise $\sigma_{\ell,i}=\sigma=\operatorname{sigmoid(\frac{z}{0.3})}$, but other options (e.g.
 rescaled softsign or Normal CFD) lead to very similar results - as long as the general sigmoid shape of the global $\sigma$ is preserved, see Section 6.2.
-Because $\sigma_{\ell}$ can be understood as measuring the excitation of a neuron we may accordingly call $\Gamma$ the excitation function, it’s corresponding pullback $v_{\Gamma}$ the excitation pullback, and it’s induced tensor field $\tilde{\Gamma}$ the excitation tensor field.
-Note that in order to compute $\Gamma$ we still need to use the original gates $G$ in the forward pass to compute the right pre-activations.
+Because $\sigma_{\ell,i}$ can be understood as measuring the excitation of a neuron we accordingly call $\Gamma$ the excitation function, it’s corresponding pullback $v_{\Gamma}$ the excitation pullback, and it’s induced tensor field $\tilde{\Gamma}$ the excitation tensor field.
+Note that in order to compute $\Gamma$ we still need to use the original gates $G$ in the forward pass to find the right pre-activations.
 Note 4.2 The tensor field $\tilde{\Gamma}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ (see Equation 12) suppresses barely active paths, i.e.
@@
 those interrupted by a single weak neuron, but otherwise strong).
-This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and probably more aligned with the coarser structure of the decision boundary.
+This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and more likely to be better aligned with the coarser structure of the decision boundary.
 In a sense, $v_{\Gamma}$ is dominated by atoms corresponding to highly excited paths, as it quickly suppresses the less-excited paths (see Equation 10).
 5 On the path stability and its potential significance
+This section should be regarded as exploratory.
+Our aim here is not to establish formal results, but rather to outline conjectures and highlight promising research avenues.
+To this end, we don’t shy away from semi-formal reasoning and intuitive arguments, leaving rigorous analysis for future work.
 5.1 The path stability hypothesis
-Because of the relative stability of $\phi_{\tilde{G}^{(t)}}$ (Note 3.3), it’s tempting to assume that the gating function $G^{(t)}$ gets fixed early on during training.
-We consider this assumption too optimistic as the gates probably do change simply because of the gradient noise we’ve hinted at before.
-Notice, however, that $\tilde{G}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ maps inputs to binary rank-1 tensors, meaning that it has limited path-filtering capability (Note 3.1).
-But perhaps there are more refined atom filters (corresponding to higher-rank binary tensors) that do get fixed early on during training?
-We motivate this assumption below.
 Earlier work has demonstrated that only a small subnetwork within a larger model is responsible for most of its performance.
@@
 Together, these observations motivate the idea that the gating patterns in ReLU networks may approximately stabilize early in training, and thus may carry semantically meaningful structure long before convergence.
-Recall that $f^{(t)}(x)=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle$ and $f^{(t)}_{\tau}(x)=\langle\omega^{(t)},\phi_{\tau}(x)\rangle$.
-Let $X=(x_{i})_{i}\subset\mathbb{R}^{d_{0}}$ be the training dataset and $f(X)=(f(x_{i}))_{i}$ be the vector of all the values of $f$.
-We say that two vectors $a,b\in\mathbb{R}^{n}$ are strongly positively correlated if the Pearson correlation coefficient $\rho(a,b)$ is high (e.g., greater than 0.9, although the exact threshold value is not important).
-We denote this relation as $a\approx_{+}b$.
+That being said we consider it too naive to assume that the gating tensors $\tilde{G}^{(t)}$ get fixed, because of their limited path-filtering capability (Note 3.1) as rank-1 binary tensors and the gradient noise we’ve hinted at before.
+However, paths may stabilize as higher-rank tensors.
+Let $\omega^{(t)}=\Omega((y^{(t)},\theta^{(t)}))\in\mathbb{R}^{P_{0}}$ encode the network’s parameters at time $t$. We’ve already established the following:
+$f^{(t)}_{\tau}(x)$ $=\langle\omega^{(t)},\phi_{\tau}(x)\rangle=\langle\omega^{(t)},\tau(x)\otimes x\rangle$ (22)
+$f^{(t)}(x)$ $=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle=\langle\omega^{(t)},\tilde{G}^{(t)}(x)\otimes x\rangle\quad\text{(substituting $\tilde{G}^{(t)}(x)$ for $\tau$)}$ (23)
+Let $X=(x_{i})_{i}\subset\mathbb{R}^{d_{0}}$ be the training dataset and $f_{\tau}(X)=(f_{\tau}(x_{i}))_{i}$ be the vector of all the values of $f_{\tau}$.
+We introduce three notions of similarity between vectors.
+We’ll say that the score is high if it exceeds some fixed value in $[0.5,1]$, understanding that the exact threshold is context-dependent.
+For $a,b\in\mathbb{R}^{n}$, we say that they are strongly positively correlated, denoted $a\sim_{corr}b$, if the Pearson correlation coefficient $\rho(a,b)$ is high.
+We say that $a$ and $b$ are approximately proportional, denoted $a\sim_{prop}b$, if there exists $\mu>0$ such that $a\sim\mu b$.
+Finally, for binary vectors $a,b\in\{0,1\}^{n}$, we say that they are consistent, denoted $a\sim_{F_{1}}b$, if their $F_{1}$-score is high, where $F_{1}(a,b)=\frac{2\sum_{i=1}^{n}a_{i}b_{i}}{\sum_{i=1}^{n}a_{i}+\sum_{i=1}^{n}b_{i}}$ is a standard measure of similarity for binary vectors.
 We now state our core hypothesis:
-Hypothesis 1 (Early emergence of stable paths).
-There exists a piecewise constant binary tensor field $\pi:\mathbb{R}^{d_{0}}\to\{0,1\}^{{P_{1}}}$ and an early training time $t_{\pi}\ll T$ such that for every $t>t_{\pi}$ we have $f^{(t)}(X)\approx_{+}f^{(t)}_{\pi}(X)$.
-Additionally, we postulate that the network encodes the tensor field $\pi$ in its pre-activation function $Z$.
-That is, there exists a choice of $(\sigma_{\ell})_{\ell}$ (as in Equation 21) such that for every $x\in X$ we have $\tilde{\Gamma}(x)\approx_{+}\pi(x)$.
-We refer to the above assumption as the path stability condition and to any such $\pi$ as path tensor feature.
-Intuition:
-The hypothesis states that, even though the gating tensors $\tilde{G}^{(t)}(x)$ change during training, there exists a time-independent gating tensor $\pi(x)$ that, starting from time $t_{\pi}$, largely determines network’s decision $f^{(t)}(x)$ at $x$.
-In this view, the variability of tensors $\tilde{G}^{(t)}(x)$ can be attributed to their limited expressive power as rank-1 tensors; the network, trying to approximate $\pi(x)$, learns it implicitly in it’s pre-activation pattern $Z(x)$ in the form of the set of highly-excited paths.
-The latter can be feasibly directionally approximated by the rank-1 non-binary tensor $\tilde{\Gamma}(x)$.
+Hypothesis 1 (Early stabilization of paths).
+There exists a piecewise constant binary tensor field $\pi:\mathbb{R}^{d_{0}}\to\{0,1\}^{{P_{1}}}$ and an early training time $t_{\pi}\ll T$ such that $\forall t>t_{\pi},\;f^{(t)}(X)\sim_{corr}f^{(t)}_{\pi}(X)\;\wedge\;\forall x\in X,\;\tilde{G}^{(t)}(x)\sim_{F_{1}}\pi(x)$.
+Additionally, tensor field $\pi$ can be chosen so that there exists a function $\Gamma$, as in Equation 21, satisfying $\forall_{x\in X}~\tilde{\Gamma}(x)\sim_{prop}\pi(x)$.
+We refer to the above assumption as the path stability condition and to any such $\pi$ as a path tensor feature.
+Motivation:
+The hypothesis formalizes the intuition that, even though the gating tensors $\tilde{G}^{(t)}(x)$ do change during training, for most $x\in X$ they actually oscillate around a core gating tensor $\pi(x)$ which determines the network’s decision $f^{(t)}(x)$ for $t>t_{\pi}$.
+In this view, the variability of the tensors $\tilde{G}^{(t)}(x)$ can be attributed to their limited expressive power as rank-1 binary approximations of $\pi(x)$.
+Additionally, we claim that $\pi(x)$ can be largely recovered as the set of those paths $p$ that are relatively highly excited, as measured by the coordinates $\tilde{\Gamma}_{p}(x)$ of the tensor $\tilde{\Gamma}(x)\in[0,1]^{\mathcal{P}_{1}}$.
+Note 5.1 The path tensor feature always exists if $t_{\pi}$ is allowed to be close to $T$, the gist of our assumption being that it forms long before convergence.
 5.2 The implicit kernel machine and the faithfulness of excitation pullbacks
-Recall that $\omega^{(t)}=\Omega((y^{(t)},\theta^{(t)}))\in\mathbb{R}^{P_{0}}$ (Equation 16) encodes the network’s parameters at time $t$ and that:
-$f^{(t)}(x)$ $=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle$
-$f^{(t)}_{\pi}(x)$ $=\langle\omega^{(t)},\phi_{\pi}(x)\rangle$
-If the Hypothesis 1 holds, then, starting from time $t_{\pi}$ the training dynamics of $\omega^{(t)}$ closely resembles the training dynamics of the kernel machine given by $f^{(t)}_{\pi}$.
-In fact, $(\phi_{\tilde{G}^{(t)}}-\phi_{\pi})$ can be interpreted as a feature noise added to the fixed feature map $\phi_{\pi}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ at time $t$ of training.
-Note 5.1 Hypothesis 1 implies that gradient descent actually learns the implicit kernel machine $f_{\pi}$ and the model $M$ is just a noisy wrapper that allows us to feasibly approximate the extremely-high-dimensional feature function $\phi_{\pi}$.
-Since $\pi$ is assumed to be piecewise constant, then also it’s pullback $v_{\pi}$ is piecewise constant and, by Equation 15, we have
+Whenever the Hypothesis 1 holds, then, starting from time $t_{\pi}$, the network becomes highly positively correlated with the kernel machine $f^{(t)}_{\pi}$ that operates on the same feature space and acts via the same vector of weights $\omega^{(t)}\in\mathbb{R}^{\mathcal{P}_{0}}$ as the network.
+In fact, $(\phi_{\tilde{G}^{(t)}}-\phi_{\pi})(x)=(\tilde{G}^{(t)}(x)-\pi(x))\otimes x$ can be interpreted as a feature noise added to the fixed feature map $\phi_{\pi}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{0}}$ at time $t$ of training.
+Thus, $M$ can itself be considered a noisy kernel machine.
+This offers an interesting angle to study the training dynamics of neural networks.
+Additionally, since $\pi$ is assumed to be piecewise constant, then so is it’s pullback vector field $v_{\pi}$ and, by Equation 15:
 $\nabla f_{\pi}(x)=v_{\pi}^{\top}(x)|_{\mathbb{R}^{d_{0}^{*}}}$ (24)
-This means that the pullback $v_{\pi}^{\top}$ can actually be interpreted as the gradient of the kernel machine $f_{\pi}$.
-We can feasibly approximate the direction of that gradient by $v_{\Gamma}$, which is a better approximation than that of $v_{G}$ and therefore $v_{\Gamma}^{\top}$ is actually more faithful to the network’s decision than the standard gradient (if the Hypothesis 1 holds).
-Also notice that in order to compute the path space dot product $\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle$ it suffices to compute the following product kernel:
+This means that $v_{\pi}^{\top}$ can be interpreted as the gradient of the kernel machine $f_{\pi}$.
+Since we have $\tilde{\Gamma}(x)\sim_{prop}\pi(x)$ by Hypothesis 1, then also $v_{\Gamma}\sim_{prop}v_{\pi}$ by Equation 14, meaning that we can directionally approximate that gradient with excitation pullback, i.e.
+the kernel machine $f_{\pi}$ is directly computable.
+Additionally, the covector $v_{\Gamma}^{\top}$ can be considered as even more faithful to the network’s decision than the standard gradient, establishing a reliable explanation method (depending on the degree to which the path stability condition is satisfied).
+Notice that, in order to find the path space dot product $\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle$, we can compute the following product kernel:
 $K_{\Lambda}(x,x^{\prime}):=\langle x,x^{\prime}\rangle\prod_{\ell=1}^{L}\langle\lambda_{\ell}(x),\lambda_{\ell}(x^{\prime})\rangle=\langle\phi_{\Lambda}(x),\phi_{\Lambda}(x^{\prime})\rangle,\quad\text{for }x,x^{\prime}\in\mathbb{R}^{d_{0}}$ (25)
-In particular, we may feasibly study the properties of $\langle\phi_{\pi}(x),\phi_{\pi}(x^{\prime})\rangle$ with the help of $K_{\Gamma}$ kernel.
-Notice that $K_{\Gamma}$ encodes the idea behind Deep Features and thus offers a possible explanation for why the latter is an effective perceptual metric (Zhang et al., 2018), as it basically expresses the path-space dot product under the feature map $\phi_{\pi}$ (up to the scaling), tapping into the internal representation of $f_{\pi}$.
+In particular, $K_{\Gamma}(x,x^{\prime})=\langle\phi_{\Gamma}(x),\phi_{\Gamma}(x^{\prime})\rangle=\langle\tilde{\Gamma}(x)\otimes x,\tilde{\Gamma}(x^{\prime})\otimes x^{\prime}\rangle\approx\langle\mu\pi(x)\otimes x,\mu^{\prime}\pi(x^{\prime})\otimes x^{\prime}\rangle=\mu\mu^{\prime}\langle\phi_{\pi}(x),\phi_{\pi}(x^{\prime})\rangle$.
+Assuming $\mu,\mu^{\prime}\approx 1$ this lets us tap into the path space dot product under the feature map $\phi_{\pi}$, positioning $K_{\Gamma}$ as a strong candidate for the measure of similarity.
+Incidentally, $K_{\Gamma}$ is equivalent to Deep Features, which is already empirically confirmed to be an effective perceptual metric (Zhang et al., 2018).
 5.3 The effectiveness of Batch Normalization
-Batch Normalization (Ioffe and Szegedy, 2015) has become one of the most influential architectural innovations in modern deep learning, largely due to its ability to dramatically accelerate training and improve generalization.
+Batch Normalization (BN) (Ioffe and Szegedy, 2015) has become one of the most influential architectural innovations in modern deep learning, largely due to its ability to dramatically accelerate training and improve generalization.
 While the standard explanation attributes these benefits to reduced internal covariate shift and smoother optimization landscapes (Santurkar et al., 2019), our framework offers an alternative and complementary perspective.
-Note that BN basically removes the bias from inputs and largely prevents the emergence of data-dominant output neurons, enforcing their uniform activation across the batch.
-This facilitates both the neuron specialization and more uniform usage of the layer width.
-Highly activated neurons are therefore considerably better aligned (directionally) with the particular input then the less activated ones and have larger impact on the weight update during training, which reinforces their role.
-All these phenomena favor quicker stabilization of path activations, i.e.
-faster emergence of the path tensor feature and the corresponding linear structure of network representations, improving the training speed and representational quality.
-Interestingly, this resonates with the recent line of work on B-cos networks (Böhle et al., 2022), which explicitly promote weight-input alignment as a mechanism for interpretability.
-The alignment of excitation pullbacks suggests that the role of B-cos transform is already performed by the BN layer in standard architectures, but this is obfuscated by the pullbacks of weakly excited neurons contaminating the gradient, see Section 4.1.
-5.4 ReLU network as content-addressable storage
-Under the path stability hypothesis, ReLU networks can be interpreted as an associative memory system with excitation pullbacks as its content.
-Recall the Note 3.1 where we interpret the pullback vector field $v_{G}$ as binary filtering of atoms.
-This can be understood as computing the target location $\tilde{G}(x)\in\mathbb{R}^{\mathcal{P}_{1}}$ in the network’s memory space $\mathbb{R}^{\mathcal{P}_{1}}$.
-During training, for a positive example $x$, the network updates its weights to pull $v_{G}^{(t)}(x)$ closer to $x$ (due to Equation 7), i.e.
-it combines $x$ with the value $v_{G^{(t)}}(x)$ stored at the location $\tilde{G}^{(t)}(x)$.
-Under the path stability condition this location remains largely the same, oscillating around $\pi(x)$ and therefore $v_{\pi}(x)$ (and $v_{\Gamma}(x)$) contains the most coherent information.
-This is validated by the perceptual alignment of excitation pullbacks.
-During inference, the network first computes the location $\tilde{G}(x)$ and then takes the dot product of $x$ with the stored memories $v_{(c,G)}(x)$ for every class $c$, returning the class of the best-matching memory.
-According to the path stability hypothesis, we can extract the actual memories by computing $v_{(c,\Gamma)}(x)$ instead (the excitation pullback for class $c$).
-It may however change the network’s prediction if the pullbacks weren’t used during the training.
+We even suspect that BN might be crucial for the early emergence of the path tensor feature.
+We sketch the arguments below, assuming that the affine parameters of BN are close to their default values (i.e.
+at the initialization or with the affine=False flag).
+The normalization performed by BN nullifies the input bias and enforces neuron specialization, as any specific unit after BN gets highly excited only for the fraction of inputs it is most aligned with (by dot product).
+Additionally, it forces every individual neuron to be inactive for around half of the inputs.
+Assuming low intra-layer correlation of weights, this encourages the network to activate a wide range of paths across the batch and only a relatively small set of paths per input, meaning that the tensors $\{\tilde{G}^{(t)}(x):x\in X\}\subset\{0,1\}^{\mathcal{P}_{1}}$ are sparse and, together, they cover a wide range of coordinates.
+This property is preserved by the tensor product and therefore transferred to the image $\phi_{\tilde{G}^{(t)}}(X)=\{\phi_{\tilde{G}^{(t)}}(x):x\in X\}\subset\mathbb{R}^{\mathcal{P}_{0}}$ under the network feature function $\phi_{\tilde{G}^{(t)}}$ (see Equation 23).
+This facilitates the linear separability of $\phi_{\tilde{G}^{(t)}}(X)$, and, in general, should significantly benefit convergence.
+Since we have not invoked Hypothesis 1 here, this reasoning can be viewed as supporting evidence in its favor.
+5.4 Classification ReLU network as content-addressable storage
+Hypothesis 1 allows us to view ReLU classification network as an associative memory with excitation pullbacks as its content.
+Recall the Note 3.1 where we interpret the pullback vector field $v_{G}$ as a binary filtering of atoms.
+This can be understood as computing the target location $\tilde{G}(x)\in\mathbb{R}^{\mathcal{P}_{1}}$ in the network’s “memory space” $\mathbb{R}^{\mathcal{P}_{1}}$.
+Given a positive (belonging to our fixed class $c$) example $x$ during training, the network updates its weights to pull $v_{G}^{(t)}(x)$ closer to $x$ (in dot product, see Equation 7), i.e.
+it combines $x$ with the value $v_{G^{(t)}}(x)$ stored at the memory location $\tilde{G}^{(t)}(x)$.
+Under the path stability condition this location remains largely the same, oscillating around $\pi(x)$.
+Moreover, Batch Normalization encourages the locations to be well-separated, as argued in Section 5.3.
+Therefore, we can expect $v_{\pi}(x)$ to store coherent features extracted from the training data distribution, which seems to be validated by the perceptual alignment of excitation pullbacks $v_{\Gamma}$.
+During inference, the network first computes the memory location $\tilde{G}(x)$ and then takes the dot product of $x$ with the stored features $v_{(c,G)}(x)$ for every class $c$, returning the class of the best-matching memory.
+According to the path stability hypothesis, we can extract the actual, coherent memories by computing $v_{(c,\Gamma)}(x)$ instead (the excitation pullback for class $c$).
+This may however change the prediction, suggesting that excitation pullbacks should be used during training in place of vanilla gradients, see Section 7.1.
 6 Empirical validation
 This section describes how we visualize the excitation pullbacks.
-To this end, we focus on popular and representative classes of ReLU-based architectures pretrained on ImageNet and available via the torchvision library, such that it is straightforward to replace most or all occurrences of ReLU and MaxPool2d (see Section 6.3) layers with our alternative variants, using a single recursive function over the child modules.
-Based on these criteria, we selected ResNet50, VGG11_BN, and DenseNet121.
-Notably, determining the optimal choice of $\sigma_{\ell}$ for every neuron or layer in a given network is outside the scope of this work.
-Instead, we seek a single global set of hyperparameters that performs reasonably well across all layers of a given model - and ideally, across all selected architectures.
-Our goal is to evaluate whether the resulting excitation pullbacks indeed exhibit a substantially improved perceptual alignment compared to standard gradients.
+We focus on popular and representative classes of ReLU-based architectures pretrained on ImageNet and available via the torchvision library, such that it is straightforward to replace most or all occurrences of ReLU and MaxPool2d (see Section 6.3) layers with our alternative variants, using a single recursive function over the child modules.
+Following these criteria, we selected ResNet50, VGG11_BN, and DenseNet121, though we obtained encouraging preliminary results on a broader range of architectures, including SiLU and GELU based.
+Note that determining the optimal choice of $\sigma_{\ell,i}$ for each neuron in a given network is left outside the scope of this work.
+This is because a single global $\sigma$ performed reasonably well across all three selected architectures, indicating significant robustness of the method to the hyperparameter choice.
+Nevertheless, we expect that tuning neuron-specific $\sigma_{\ell,i}$ functions could substantially improve the quality of feature attributions.
 6.1 Experimental setup
@@
 Notice that $\Gamma$ approaches hard gating $G$ as temp goes to zero and no gating at all as temp goes to infinity.
-We use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True from the torchvision.models library.555see Section 7.3 for more context.
+We use resnet50, vgg11_bn and densenet121 models with the flag pretrained=True from the torchvision.models library (see Section 7.2).
 Gradients are visualised by torchvision.utils.make_grid(scale_each=True).
@@
 B-SiLU) consistently improves generalization performance.
-Combined with our findings this leads to the open question, whether computing excitation pullbacks (instead of vanilla gradients) during training can improve generalization, and to what extent.
+Combined with our findings this leads to the question, whether computing excitation pullbacks (instead of vanilla gradients) during training can improve both generalization and interpretability, including other activations.
 Intuitively, one should not differentiate through the excitation gates explicitly as they may saturate and overfit; nevertheless, they can still be computed in the backward pass as alternative gradients.
@@
 Furthermore, the faithfulness of the produced explanations is established intuitively, without referencing the training dynamics.
+Our method resonates with the recent line of work on B-cos networks (Böhle et al., 2022), which explicitly promotes weight-input alignment as a mechanism for interpretability.
+The perceptual alignment of excitation pullbacks suggests that the role of B-cos transform is already performed by the BN layer, which aligns highly excited neurons with inputs, but this is obfuscated by the pullbacks of weakly excited neurons contaminating the gradients (see Section 4.1).
 It has been observed that networks trained to be robust to adversarial perturbations yield gradients that are often aligned with human perception, a phenomenon first observed in (Tsipras et al., 2019) and then referred to as perceptually-aligned gradients (PAG) in (Kaur et al., 2019).
@@
 This resonates with our observation that excitation pullback $v_{\Gamma}$ provides a smoother approximation of the model’s decision boundary then the standard, hard-gated gradients and therefore, by extension, may better align with the actual data manifold.
-7.2 On the possible biological plausibility of the implicit linear model
-Intuitively, the relative stability of the gating mechanism throughout training is reminiscent of the approximately fixed synaptic connections observed in biological neural systems.
-In this analogy, the discrete gating pattern can be thought of as defining a persistent connectivity structure - akin to neural connections - while the excitation function describes the level of activity (excitation) of each neuron, resembling the frequency-based coding of information.
-One may even venture a more speculative interpretation: the excitation pullbacks could be loosely associated with internal subjective representations, i.e.
-qualia - subjective, localized manifestations of experience that emerge from the brain activity taken as a whole - in a way similar to how excitation pullbacks emerge from the gating patterns in ReLU networks.
-It makes even more sense in the light of Section 5.4, where we interpret excitation pullbacks as the content of the network’s associative memory.
-This may open exciting direction in neuroscience and the philosophy of mind.
 7.2 Limitations
@@
 This flag is now deprecated in the recent versions of the library, in particular it returns the older ResNet50_Weights.IMAGENET1K_V1.
-We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights, at least when using a single global $\sigma$ for all the neurons, which, as we’ve indicated earlier, might by an oversimplified approach.
-We speculate that this is because of different distribution of pre-activations in newer versions of the pretrained models, which have been heavily regularized.
+We’ve observed that it’s harder (but still possible) to achieve similar quality of excitation pullbacks for the more recent ResNet50_Weights.IMAGENET1K_V2 weights.
+We speculate that this is because of different distribution of pre-activations in newer, heavily regularized versions of pretrained models.
 The formal theory developed in this paper concerns specifically the ReLU networks.
@@
 SiLU and GELU.
-Initial experiments for those architectures show promise but require certain refinement of the computation of $\Gamma$ to account for the more complex gradients.
+Initial experiments for those architectures show promise but require modifying $\Gamma$ to account for the more complex gradients.
 We haven’t yet explored the potential applicability of our method to attention-based models.
@@
 This underlines the importance of an appropriate inductive bias for our results to carry over to other domains.
+A key limitation of our work is that the contributions remain preliminary and partly speculative.
+A more rigorous formal treatment and supporting empirical studies are required, which we defer to future work.
 8 Acknowledgments